### v2

# Well-posedness of Hibler’s parabolic-hyperbolic sea ice model

Abstract.
This paper proves the local-in-time strong well-posedness of a parabolic-hyperbolic regularized version of Hibler’s sea ice model. Hibler’s model is the most frequently used sea ice model in climate science. Lagrangian coordinates are employed to handle the hyperbolic terms in the balance laws. The resulting problem is regarded as a quasilinear non-autonomous evolution equation. Maximal $\mathrm{L}^{p}$-regularity of the underlying linearized problem is obtained on an anisotropic ground space in order to deal with the lack of regularization in the balance laws.
Key words and phrases:
Hibler’s sea ice model, local strong well-posedness, viscous-plastic material, parabolic-hyperbolic problem, quasilinear evolution equation, Lagrangian coordinates, anisotropic ground space
2020 Mathematics Subject Classification:
35Q86, 35K59, 86A05, 86A10
1. Introduction
Recent years have witnessed an increase in research on sea ice dynamics both, from the theoretical and the applied side. One reason for this might be the role of sea ice in climate science, acting as an insulator in between the atmosphere and the ocean. For more details on the importance of sea ice in climate studies, we refer e. g. to the survey article of Hunke, Lipscomb and Turner [15]. Golden et al. [12] provide an overview of sea ice modeling. To date, almost all climate models capturing sea ice dynamics rely on Hibler’s viscous-plastic model introduced in [13] in 1979. It is a 2D large-scale model that describes dynamic as well as thermodynamic aspects of sea ice. In particular, the complex mechanical behavior of sea ice is taken into account in the viscous-plastic rheology.
First well-posedness results to Hibler’s model were obtained by Brandt, Disser, Haller-Dintelmann and Hieber [6] as well as Liu, Thomas and Titi [23]. In [6], a fully parabolic variant of Hibler’s model is investigated, and local-in-time strong well-posedness and global-in-time well-posedness close to equilibrium solutions are shown. The parabolic-hyperbolic nature of the model is stressed in [23], but the authors strongly modify the stress tensor and establish local strong well-posedness of the resulting problem.
It is the aim of the present paper to prove the local strong well-posedness of the parabolic-hyperbolic variant of Hibler’s model, where the most commonly used regularization for the stress tensor is employed. Unlike the models under consideration in [6] or [23], this is precisely the model typically studied in numerical simulations.
In contrast to the rigorous mathematical analysis, which only started quite recently, there is vast literature on numerical analysis and simulations of Hibler’s model, see for example the work of Zhang and Hibler [37]. In order to ease computational costs, Hunke and Dukowicz [14] suggested an elastic-viscous-plastic rheology. This model was later on modified in order to reduce the discrepancy with observations, see e. g. the article of Kimmrich, Danilov and Losch [16]. Lemieux and Tremblay [20] implemented an inexact Newton method for Hibler’s model. In [26], Mehlmann and Richter developed a modified Newton solver. For further works in this direction, we also refer to the articles of Seinen and Khouider [33], Shih, Mehlmann, Losch and Stadler [34] or Yaremchuk and Panteleev [36]. Mehlmann and Korn [25] provided numerical studies of sea ice on triangular grids for an integration into the ocean general circulation model ICON-O.
As indicated above, we invoke a regularized stress tensor here. In fact, although the original stress tensor describes an idealized viscous-plastic material, it gives rise to a degenerate problem, see also Section 2. As a result, already Hibler [13] suggested to cut off the viscosities. In the present paper, we follow the most common regularization of the stress tensor as e. g. suggested by Kreyscher et al. [18]. Recently, different possibilities to modify the rheology were suggested by Chatta, Khouider and Kesri [8] and Ringeisen, Losch and Tremblay [32].
For the analysis, we employ the Lagrangian change of coordinates in order to deal with the hyperbolic effects in the balance laws. The advantage is that the hyperbolic terms in the balance laws are annihilated by the introduction of Lagrangian coordinates. More precisely, the time derivative with respect to Lagrangian coordinates corresponds precisely to the material derivative in the Eulerian formulation. For the Lagrangian approach in the context of the compressible Navier-Stokes equations, we refer for instance to the article of Danchin [9]. Let us stress, however, that the present problem differs significantly from compressible Navier-Stokes equations. First, the regularized stress tensor $\sigma_{\delta}$ shown in (2.2) depends on the deformation tensor $\varepsilon=\nicefrac{{1}}{{2}}(\nabla_{\mathrm{H}}v_{\mathrm{ice}}+(\nabla_{\mathrm{H}}v_{\mathrm{ice}})^{\top})$ in a nonlinear way. In addition, the pressure is replaced by a so-called ice strength $P=P(h,a)$ as displayed in (2.1) which is in turn a function of the ice thickness $h$ and the ice concentration $a$ for which balance laws are prescribed.
For the principal variable $u=(v_{\mathrm{ice}},h,a)$ capturing the horizontal ice velocity $v_{\mathrm{ice}}$, the ice thickness $h$ and the ice concentration $a$, a suitable operator $A(t,u)$ and right-hand side $F(t,u)$, we rewrite the resulting non-autonomous and quasilinear system in Lagrangian coordinates as a Cauchy problem of the form
$\frac{\mathrm{d}}{\mathrm{d}t}u+A(t,u)u=F(t,u),\enspace\text{for}\enspace t>0,\enspace u(0)=u_{0},$
on a Banach space $\mathrm{X}_{0}$. For the linearized problem, we employ the so-called maximal regularity method. Roughly speaking, this means that for $A(u_{0})\coloneqq A(0,u_{0})\colon\mathrm{X}_{1}\to\mathrm{X}_{0}$ and the time trace $\operatorname{tr}$, the operator
$\Bigl{(}\frac{\mathrm{d}}{\mathrm{d}t}+A(u_{0}),\operatorname{tr}\Bigr{)}\colon\mathbb{E}_{1}\to\mathbb{E}_{0}\times\mathrm{X}_{\gamma}$
is an isomorphism between suitable Banach spaces $\mathbb{E}_{1}$ and $\mathbb{E}_{0}\times\mathrm{X}_{\gamma}$. In this paper, we choose the $\mathrm{L}^{p}$-framework, so we consider solutions in $\mathbb{E}_{1}\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1})$, meaning that every term in the differential equation lies in $\mathbb{E}_{0}\coloneqq\mathrm{L}^{p}(0,T;\mathrm{X}_{0})$. The space $\mathrm{X}_{\gamma}$ then is the real interpolation space $(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$. For further details on maximal regularity in general and maximal $\mathrm{L}^{p}$-regularity in particular, we refer for instance to the monographs of Amann [2], Lunardi [24] or Prüss and Simonett [30], the memoir of Denk, Hieber and Prüss [11] or the book chapter of Kunstmann and Weis [19]. It is a key step in our analysis to introduce an anisotropic ground space of the form $\mathrm{X}_{0}=\mathrm{L}^{q}\times\mathrm{W}^{1,q}\times\mathrm{W}^{1,q}$ for the ice velocity, ice thickness and ice concentration to obtain the maximal $\mathrm{L}^{p}$-regularity of the linearized problem. In fact, we even show that the associated operator matrix admits a bounded $\mathcal{H}^{\infty}$-calculus by means of theory for diagonally dominant block operator matrices and deduce the maximal $\mathrm{L}^{p}$-regularity from there. We then use the maximal $\mathrm{L}^{p}$-regularity to set up a fixed point argument in the Lagrangian formulation. For this, a precise tracking of the change of coordinates is required to derive suitable nonlinear estimates.
Recently, sea ice has also been studied in different contexts. In [7], sea ice is investigated in the time periodic setting, see also [5, Section 7.2]. Moreover, local strong well-posedness of the interaction problem of sea ice with a rigid body is established in [3]. The paper [4] introduces sea ice as a thin layer in between the atmosphere and the ocean and establishes local strong well-posedness as well as global strong well-posedness close to equilibria of the coupled model. The motivation is the coupled atmosphere-ocean model introduced by Lions, Temam and Wang [21, 22]. In the papers [7, 3, 4], the fully parabolic regularized variant of Hibler’s model as analyzed in [6] is taken into account. Let us also mention the recent work of Piersanti and Temam [29] on the modeling and analysis of the dynamics of shallow ice sheets.
This article is organized as follows. In Section 2, we introduce Hibler’s model and state the main result, Theorem 2.1, on the local-in-time strong well-posedness. Section 3 is then dedicated to the transformation to Lagrangian coordinates and to the reformulation of the main result, see Theorem 3.1. In addition, we discuss the continuous dependence on the initial data as well as a blow-up criterion in the Lagrangian formulation. The focal point of Section 4 is the proof of the bounded $\mathcal{H}^{\infty}$-calculus and the maximal $\mathrm{L}^{p}$-regularity of the linearized operator matrix with the anisotropic ground space. In Section 5, we prove the main result. For this, we start with the reformulation as a fixed point problem. Thereafter, we derive the estimates of the nonlinear terms. We then deduce Theorem 2.1 from Theorem 3.1 by the invertibility of the change of coordinates for small time. Section 6 presents some concluding remarks.
2. Hibler’s sea ice model and main result
In this section, we first recall Hibler’s viscous-plastic sea ice model from [13] and then state the main result on the local strong well-posedness. In the sequel, we consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with boundary $\partial\Omega$ of class $\mathrm{C}^{2}$ and a time interval $(0,T)$, where $0<T<\infty$. The model variables are the horizontal sea ice velocity $v_{\mathrm{ice}}\colon(0,T)\times\Omega\to\mathbb{R}^{2}$, the mean ice thickness $h\colon(0,T)\times\Omega\to[\kappa,\infty)$, with $\kappa>0$ denoting a small parameter, and the ice concentration $a\colon(0,T)\times\Omega\to(0,1)$. The latter variable describes the ratio of area covered by thick ice, i. e., sea ice with a thickness larger than $h_{\bullet}$, where $h_{\bullet}>0$ is a certain thickness level. The lower bound of the mean ice thickness means that in each control area, there is at least some sea ice. In the sequel, we denote by $u$ the principle variable, i. e., $u=(v_{\mathrm{ice}},h,a)$.
The characteristics of sea ice as a viscous-plastic material are encoded in the stress tensor introduced below. In the following, $\varepsilon=\varepsilon(v_{\mathrm{ice}})=\nicefrac{{1}}{{2}}(\nabla_{\mathrm{H}}v_{\mathrm{ice}}+(\nabla_{\mathrm{H}}v_{\mathrm{ice}})^{\top})$ represents the deformation tensor. We remark that $\nabla_{\mathrm{H}}$ is the gradient in two variables. The same is valid for $\mathrm{div}_{\mathrm{H}}\,$, denoting the horizontal divergence. Moreover, for given constants $p^{*}>0$ and $c_{\bullet}>0$, the ice strength $P$ takes the explicit shape
(2.1) $P=P(h,a)=p^{*}h\mathrm{e}^{-c_{\bullet}(1-a)}.$
For the ratio $e>1$ of the major to minor axes of the elliptical yield curve of the principle components of the stress, we introduce
$\triangle^{2}(\varepsilon)\coloneqq\left(\varepsilon_{11}^{2}+\varepsilon_{22}^{2}\right)\Bigl{(}1+\frac{1}{e^{2}}\Bigr{)}+\frac{4}{e^{2}}\varepsilon_{12}^{2}+2\varepsilon_{11}\varepsilon_{22}\Bigl{(}1-\frac{1}{e^{2}}\Bigr{)}.$
The resulting bulk and shear viscosities then take the shape
$\zeta(\varepsilon,P)=\frac{P(h,a)}{2\triangle(\varepsilon)}\enspace\text{and}\enspace\eta(\varepsilon,P)=e^{-2}\zeta(\varepsilon,P).$
The stress tensor $\sigma$ accounting for the behavior of sea ice as a viscous-plastic material then reads as
$\sigma=2\eta(\varepsilon,P)\varepsilon+[\zeta(\varepsilon,P)-\eta(\varepsilon,P)]\operatorname{tr}(\varepsilon)\operatorname{Id}_{2}-\frac{P}{2}\operatorname{Id}_{2},$
with $\operatorname{Id}_{2}$ denoting the identity matrix. In spite of the fact that the above law describes an idealized viscous-plastic behavior, the bulk and shear viscosities $\zeta$ and $\eta$ become singular for $\varepsilon\to 0$. For this reason, we invoke the same regularization as in [25] and [6], see also [18], by defining
$\triangle_{\delta}(\varepsilon)\coloneqq\sqrt{\delta+\triangle^{2}(\varepsilon)}$
for $\delta>0$. Accordingly, we introduce the regularized viscosities
$\zeta_{\delta}(\varepsilon,P)\coloneqq\frac{P(h,a)}{2\triangle_{\delta}(\varepsilon)}\enspace\text{and}\enspace\eta_{\delta}(\varepsilon,P)\coloneqq e^{-2}\zeta_{\delta}(\varepsilon,P),$
resulting in the regularized stress tensor $\sigma_{\delta}$ studied in the remainder of this article and given by
(2.2) $\sigma_{\delta}\coloneqq 2\eta_{\delta}(\varepsilon,P)\varepsilon+[\zeta_{\delta}(\varepsilon,P)-\eta_{\delta}(\varepsilon,P)]\operatorname{tr}(\varepsilon)\operatorname{Id}_{2}-\frac{P}{2}\operatorname{Id}_{2}.$
Hibler’s model consists of a momentum equation for the sea ice velocity $v_{\mathrm{ice}}$ coupled to balance laws for $h$ and $a$. The momentum balance includes the internal ice stress expressed by $\mathrm{div}_{\mathrm{H}}\,\sigma_{\delta}$ as well as external forcing terms and takes the shape
$m_{\mathrm{ice}}\left(\partial_{t}v_{\mathrm{ice}}+(v_{\mathrm{ice}}\cdot\nabla_{\mathrm{H}})v_{\mathrm{ice}}\right)=\mathrm{div}_{\mathrm{H}}\,\sigma_{\delta}-m_{\mathrm{ice}}c_{\mathrm{cor}}v_{\mathrm{ice}}^{\perp}-m_{\mathrm{ice}}g\nabla_{\mathrm{H}}H+\tau_{\mathrm{atm}}+\tau_{\mathrm{ocn}}(v_{\mathrm{ice}}).$
In the above $m_{\mathrm{ice}}=\rho_{\mathrm{ice}}h$ represents the ice mass for a constant ice density $\rho_{\mathrm{ice}}>0$, $c_{\mathrm{cor}}>0$ is the Coriolis parameter, $g$ denotes the gravity, $H\colon(0,T)\times\Omega\to[0,\infty)$ is the sea surface dynamic height and $\tau_{\mathrm{atm}}$ as well as $\tau_{\mathrm{ocn}}(v_{\mathrm{ice}})$ correspond to the atmospheric wind and oceanic forces. The latter two read as
$\tau_{\mathrm{atm}}=\rho_{\mathrm{atm}}C_{\mathrm{atm}}|V_{\mathrm{atm}}|R_{\mathrm{atm}}V_{\mathrm{atm}}\enspace\text{and}\enspace\tau_{\mathrm{ocn}}(v_{\mathrm{ice}})=\rho_{\mathrm{ocn}}C_{\mathrm{ocn}}|V_{\mathrm{ocn}}-v_{\mathrm{ice}}|R_{\mathrm{ocn}}(V_{\mathrm{ocn}}-v_{\mathrm{ice}})$
for densities $\rho_{\mathrm{atm}}>0$ and $\rho_{\mathrm{ocn}}>0$, drag coefficients $C_{\mathrm{atm}}>0$ and $C_{\mathrm{ocn}}>0$, velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ of the atmosphere and the ocean and rotation matrices $R_{\mathrm{atm}}$ and $R_{\mathrm{ocn}}$.
The balance laws of the mean ice thickness $h$ and the ice concentration $a$ are given by
[…]
where $S_{\mathrm{h}}$ and $S_{\mathrm{a}}$ represent thermodynamic source terms. For a function $f_{\mathrm{gr}}\in\mathrm{C}_{\mathrm{b}}^{1}([0,\infty))$ modeling the ice growth rate, see for instance the one suggested by Hibler [13], these terms are
$S_{\mathrm{h}}(h,a)$ $=f_{\mathrm{gr}}\left(\frac{h}{a}\right)a+(1-a)f_{\mathrm{gr}}(0)\enspace\text{and}\enspace$
[…]
In the present setting, the model is completed by Dirichlet boundary conditions for the sea ice velocity, so $v_{\mathrm{ice}}=0$ on $(0,T)\times\partial\Omega$, while no boundary conditions for $h$ and $a$ are assumed. Moreover, the initial conditions read as $v_{\mathrm{ice}}(0)=v_{\mathrm{ice},0}$, $h(0)=h_{0}$ and $a(0)=a_{0}$.
The previous considerations are summarized in the complete system of equations
(2.3) $\partial_{t}v_{\mathrm{ice}}+(v_{\mathrm{ice}}\cdot\nabla_{\mathrm{H}})v_{\mathrm{ice}}&=\frac{1}{m_{\mathrm{ice}}}\mathrm{div}_{\mathrm{H}}\,\sigma_{\delta}-c_{\mathrm{cor}}v_{\mathrm{ice}}^{\perp}-g\nabla_{\mathrm{H}}H\\ &\quad+\frac{1}{m_{\mathrm{ice}}}(\tau_{\mathrm{atm}}+\tau_{\mathrm{ocn}}(v_{\mathrm{ice}})),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}h+\mathrm{div}_{\mathrm{H}}\,(v_{\mathrm{ice}}h)&=S_{\mathrm{h}}(h,a),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}a+\mathrm{div}_{\mathrm{H}}\,(v_{\mathrm{ice}}a)&=S_{\mathrm{a}}(h,a),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ v_{\mathrm{ice}}&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\Omega,\\ v_{\mathrm{ice}}(0)=v_{\mathrm{ice},0},\enspace h(0)&=h_{0},\enspace a(0)=a_{0},&&\enspace\text{in}\enspace\Omega.\end{aligned}\right.$
Before stating the main result, we require some further preparation with regard to notation which we address below. For brevity, we define the spaces $\mathrm{X}_{0}$ and $\mathrm{X}_{1}$ by
(2.4) $\mathrm{X}_{0}\coloneqq\mathrm{L}^{q}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)\enspace\text{and}\enspace\mathrm{X}_{1}\coloneqq\mathrm{W}^{2,q}(\Omega)^{2}\cap\mathrm{W}_{0}^{1,q}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega),$
where $q\in(1,\infty)$. With regard to the initial values, for $p\in(1,\infty)$, we then introduce the time trace space $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$ taking the shape
(2.5) $\mathrm{X}_{\gamma}=\mathrm{B}_{qp,\mathrm{D}}^{2-\nicefrac{{2}}{{p}}}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega).$
In the above, the subscript D indicates Dirichlet boundary conditions. If $p$, $q\in(1,\infty)$ satisfy
(2.6) $\frac{1}{p}+\frac{1}{q}<\frac{1}{2},$
then it follows from Sobolev embeddings, see for example [35, Theorem 4.6.1], that
(2.7) $\mathrm{X}_{\gamma}\hookrightarrow\mathrm{B}_{qp}^{2-\nicefrac{{2}}{{p}}}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)\hookrightarrow\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{C}^{0,\alpha}(\overline{\Omega})\times\mathrm{C}^{0,\alpha}(\overline{\Omega}),$
for some $\alpha>0$. In order to guarantee that the initial values lie in the physically relevant ranges of the respective variables, we introduce an open set $V\subset\mathrm{X}_{\gamma}$ with
(2.8) $V\coloneqq\{u=(v_{\mathrm{ice}},h,a)\in\mathrm{X}_{\gamma}:h>\kappa\enspace\text{and}\enspace a\in(0,1)\}.$
In view of (2.7), it is legitimate to make pointwise assumptions on $h$ and $a$ for $u\in\mathrm{X}_{\gamma}$ if (2.6) is valid.
For the statement of the main result, we explicitly state the assumptions on the external forcing terms.
[…]
Let $q\in(1,\infty)$, and let $T_{1}>0$ sufficiently large to be determined precisely later. We make the following assumptions on the data.
  1. (a)
    The wind velocity $V_{\mathrm{atm}}$ and the ocean velocity $V_{\mathrm{ocn}}$ satisfy $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{2q}(\Omega)^{2})$.
  2. (b)
    For the sea surface dynamic height $H$, it is valid that $\nabla_{\mathrm{H}}H\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{q}(\Omega)^{2})$.
  3. (c)
    It holds that $f_{\mathrm{gr}}\in\mathrm{C}_{\mathrm{b}}^{1}([0,\infty))$ for the ice growth rate $f_{\mathrm{gr}}$.
We are now in the position to formulate the main result of this article.
Theorem 2.1.
Let $p,q\in(1,\infty)$ be such that (2.6) is valid, let $u_{0}\in V$, where $V$ was introduced in (2.8), and recall the spaces $\mathrm{X}_{0}$ and $\mathrm{X}_{1}$ from (2.4). Moreover, suppose that the external terms $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}$, $H$ and $f_{\mathrm{gr}}$ fulfill Section 2. Then there is $T>0$ such that (2.3) has a unique solution $u=(v_{\mathrm{ice}},h,a)$ with
$u\in\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1})\cap\mathrm{C}([0,T];V).$
3. Transformation to Lagrangian coordinates
This section is dedicated to the transformation to Lagrangian coordinates which is a major part of the strategy to show the local strong well-posedness. The coordinate transform from Eulerian to Lagrangian coordinates involves the characteristics associated to the ice velocity $v_{\mathrm{ice}}$. In other words, $X$ solves
(3.1) $\partial_{t}X(t,y_{\mathrm{H}})&=v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}})),&&\enspace\text{for}\enspace t>0,\\ X(0,y_{\mathrm{H}})&=y_{\mathrm{H}},&&\enspace\text{for}\enspace y_{\mathrm{H}}\in\mathbb{R}^{2}.\end{aligned}\right.$
Given $t\in(0,T)$, we use $Y(t,\cdot)=[X(t,\cdot)]^{-1}$ to denote the inverse of $X(t,\cdot)$, and we will elaborate on the invertibility in more details later on. The variables in Lagrangian coordinates are then defined by
(3.2) $\tilde{v}_{\mathrm{ice}}(t,y_{\mathrm{H}})$ $\coloneqq v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}})),\enspace\tilde{h}(t,y_{\mathrm{H}})\coloneqq h(t,X(t,y_{\mathrm{H}}))\enspace\text{and}\enspace\tilde{a}(t,y_{\mathrm{H}})\coloneqq a(t,X(t,y_{\mathrm{H}})).$
Integrating in time in (3.1), and inserting the change of variables from (3.2), we deduce that
$X(t,y_{\mathrm{H}})=y_{\mathrm{H}}+\int_{0}^{t}\tilde{v}_{\mathrm{ice}}(s,y_{\mathrm{H}})\,\mathrm{d}s.$
The main reason for the transform of the Lagrangian coordinates is that this allows us to circumvent the hyperbolic effects in the balance laws. In fact, it readily follows that the time derivative in Lagrangian coordinates coincides with the material derivative in Eulerian coordinates. Indeed, we have
$\partial_{t}\tilde{v}_{\mathrm{ice}}(t,y_{\mathrm{H}})$ $=\partial_{t}v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}}))+(v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}}))\cdot\nabla_{\mathrm{H}})v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}})),$
$\partial_{t}\tilde{h}(t,y_{\mathrm{H}})$ $=\partial_{t}h(t,X(t,y_{\mathrm{H}}))+v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}}))\cdot\nabla_{\mathrm{H}}h(t,X(t,y_{\mathrm{H}}))\enspace\text{and}\enspace$
$\partial_{t}\tilde{a}(t,y_{\mathrm{H}})$ $=\partial_{t}a(t,X(t,y_{\mathrm{H}}))+v_{\mathrm{ice}}(t,X(t,y_{\mathrm{H}}))\cdot\nabla_{\mathrm{H}}a(t,X(t,y_{\mathrm{H}})).$
We next compute the other transformed terms from (2.3). For the internal ice stress, we first recall the so-called Hibler operator related to $\mathrm{div}_{\mathrm{H}}\,\sigma_{\delta}$, and we define the matrix $\mathbb{S}\colon\mathbb{R}^{2\times 2}\to\mathbb{R}^{2\times 2}$ such that
$\mathbb{S}\varepsilon=\begin{pmatrix}\left(1+\frac{1}{e^{2}}\right)\varepsilon_{11}+\left(1-\frac{1}{e^{2}}\right)\varepsilon_{22}&\frac{1}{e^{2}}\left(\varepsilon_{12}+\varepsilon_{21}\right)\\ \frac{1}{e^{2}}\left(\varepsilon_{12}+\varepsilon_{21}\right)&\left(1-\frac{1}{e^{2}}\right)\varepsilon_{11}+\left(1+\frac{1}{e^{2}}\right)\varepsilon_{22}\end{pmatrix}.$
An identification of $\varepsilon\in\mathbb{R}^{2\times 2}$ with the vector $(\varepsilon_{11},\varepsilon_{12},\varepsilon_{21},\varepsilon_{22})^{\top}\in\mathbb{R}^{4}$ then allows us to interpret the action of $\mathbb{S}$ to $\varepsilon$ as the multiplication by the matrix
$\mathbb{S}=\left(\mathbb{S}_{ij}^{kl}\right)=\begin{pmatrix}1+\frac{1}{e^{2}}&0&0&1-\frac{1}{e^{2}}\\ 0&\frac{1}{e^{2}}&\frac{1}{e^{2}}&0\\ 0&\frac{1}{e^{2}}&\frac{1}{e^{2}}&0\\ 1-\frac{1}{e^{2}}&0&0&1+\frac{1}{e^{2}}\end{pmatrix}.$
This also leads to $\triangle^{2}(\varepsilon)=\varepsilon^{\top}\mathbb{S}\varepsilon$. Defining $S_{\delta}\coloneqq S_{\delta}(\varepsilon,P)\coloneqq\frac{P}{2}\frac{\mathbb{S}\varepsilon}{\triangle_{\delta}(\varepsilon)}$, we find $\sigma_{\delta}(\varepsilon,P)=S_{\delta}(\varepsilon,P)-\frac{P}{2}\operatorname{Id}_{2}$. Hibler’s operator is then defined by
$\mathbb{A}^{\mathrm{H}}v_{\mathrm{ice}}\coloneqq\frac{1}{\rho_{\mathrm{ice}}h}\mathrm{div}_{\mathrm{H}}\,S_{\delta}=\frac{1}{\rho_{\mathrm{ice}}h}\mathrm{div}_{\mathrm{H}}\,\left(\frac{P}{2}\frac{\mathbb{S}\varepsilon}{\sqrt{\delta+\varepsilon^{\top}\mathbb{S}\varepsilon}}\right).$
From [6, Section 3], we recall its representation
(3.3) $(\mathbb{A}^{\mathrm{H}}v_{\mathrm{ice}})_{i}$ $=\sum_{j,k,l=1}^{2}\frac{P}{2\rho_{\mathrm{ice}}h}\frac{1}{\triangle_{\delta}(\varepsilon)}\Bigl{(}\mathbb{S}_{ij}^{kl}-\frac{1}{\triangle_{\delta}^{2}(\varepsilon)}(\mathbb{S}\varepsilon)_{ik}(\mathbb{S}\varepsilon)_{lj}\Bigr{)}\partial_{k}\varepsilon_{jl}+\frac{1}{2\rho_{\mathrm{ice}}h\triangle_{\delta}(\varepsilon)}\sum_{j=1}^{2}(\partial_{j}P)(\mathbb{S}\varepsilon)_{ij},$
where $i=1,2$. For the principal part, it is advantageous to define coefficients by
(3.4) $a_{ij}^{kl}(\varepsilon,P)\coloneqq-\frac{P}{2\rho_{\mathrm{ice}}h}\frac{1}{\triangle_{\delta}(\varepsilon)}\left(\mathbb{S}_{ij}^{kl}-\frac{1}{\triangle_{\delta}^{2}(\varepsilon)}(\mathbb{S}\varepsilon)_{ik}(\mathbb{S}\varepsilon)_{jl}\right).$
At this stage, it is also worthwhile invoking the linearized Hibler operator. For $u_{0}\in\mathrm{C}^{1}(\overline{\Omega})^{2}\times\mathrm{C}(\overline{\Omega})\times\mathrm{C}(\overline{\Omega})$ such that $h_{0}>\kappa$, and with the notation $\mathrm{D}_{m}=-\mathrm{i}\partial_{m}$, it is of the form
(3.5) i $=\sum_{j,k,l=1}^{2}a_{ij}^{kl}(\varepsilon(v_{\mathrm{ice},0}),P(h_{0},a_{0}))\mathrm{D}_{k}\mathrm{D}_{l}v_{\mathrm{ice},j}$
$\quad+\frac{1}{2\rho_{\mathrm{ice}}h_{0}\triangle_{\delta}(\varepsilon(v_{\mathrm{ice},0}))}\sum_{j=1}^{2}(\partial_{j}P(h_{0},a_{0}))(\mathbb{S}\varepsilon(v_{\mathrm{ice}}))_{ij}.$
The above conditions on $u_{0}$ are especially satisfied provided $u_{0}\in V$, and $p,q\in(1,\infty)$ fulfill (2.6).
Concerning the transformed terms, we start with the symmetric part of the gradient, transforming to
$2\varepsilon_{ij}(v_{\mathrm{ice}})=\partial_{i}v_{\mathrm{ice},j}+\partial_{j}v_{\mathrm{ice},i}=\sum_{k=1}^{2}(\partial_{i}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}+(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},i}\eqqcolon 2\tilde{\varepsilon}_{ij}(\tilde{v}_{\mathrm{ice}}).$
For the coefficients as in (3.4), we also define
(3.6) $a_{ij}^{klm}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}),P(\tilde{h},\tilde{a}))\coloneqq(\partial_{k}Y_{m})a_{ij}^{kl}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}),P(\tilde{h},\tilde{a})).$
Besides, we compute
$\partial_{m}\tilde{\varepsilon}_{jl}(v_{\mathrm{ice}})=\frac{1}{2}\sum_{n=1}^{2}\bigl{(}(\partial_{m}\partial_{j}Y_{n})\partial_{n}\tilde{v}_{ice,l}+(\partial_{j}Y_{n})\partial_{m}\partial_{n}\tilde{v}_{\mathrm{ice},l}+(\partial_{m}\partial_{l}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}+(\partial_{l}Y_{n})\partial_{m}\partial_{n}\tilde{v}_{\mathrm{ice},j}\bigr{)}.$
In view of (3.3), these calculations result in the transformed Hibler operator given by
(3.7) $\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u})\tilde{v}_{\mathrm{ice}}$ $=\sum_{j,k,l,m=1}^{2}a_{ij}^{klm}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}),P(\tilde{h},\tilde{a}))\partial_{m}\tilde{\varepsilon}_{jl}(\tilde{v}_{\mathrm{ice}})$
$\quad+\frac{1}{2\rho_{\mathrm{ice}}\tilde{h}\triangle_{\delta}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}))}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})(\partial_{k}\tilde{h}+c_{\bullet}\partial_{k}\tilde{a})(\mathbb{S}\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}))_{ij}.$
The transformed terms associated to the horizontal divergence of the ice strength $P$ read as
$(\tilde{B}_{\mathrm{h}}(\tilde{u})\tilde{h})_{i}\coloneqq\frac{\partial_{h}P(\tilde{h},\tilde{a})}{2\rho_{\mathrm{ice}}\tilde{h}}\sum_{j=1}^{2}(\partial_{i}Y_{j})\partial_{j}\tilde{h}\enspace\text{and}\enspace(\tilde{B}_{\mathrm{a}}(\tilde{u})\tilde{a})_{i}\coloneqq\frac{\partial_{a}P(\tilde{h},\tilde{a})}{2\rho_{\mathrm{ice}}\tilde{h}}\sum_{j=1}^{2}(\partial_{i}Y_{j})\partial_{j}\tilde{a}.$
For brevity, we will also use the notation $\tilde{B}(\tilde{u})\binom{\tilde{h}}{\tilde{a}}\coloneqq\tilde{B}_{\mathrm{h}}(\tilde{u})\tilde{h}+\tilde{B}_{\mathrm{a}}(\tilde{u})\tilde{a}$ in the sequel. The last transformed term involving differential operators is $\mathrm{div}_{\mathrm{H}}\,v_{\mathrm{ice}}$, and we get $\mathrm{div}_{\mathrm{H}}\,v_{\mathrm{ice}}=\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}$. The remaining transformed terms are obtained by an insertion of the variables in Lagrangian coordinates instead of Eulerian coordinates, i. e., $u=(v_{\mathrm{ice}},h,a)$ is replaced by $(\tilde{v}_{\mathrm{ice}},\tilde{h},\tilde{a})$. In summary, with the abbreviation $\tau_{\mathrm{ice}}(\tilde{v}_{\mathrm{ice}})\coloneqq\tau_{\mathrm{atm}}+\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice}})$, the transformed system of equations takes the shape
(3.8) $\partial_{t}\tilde{v}_{\mathrm{ice}}&=\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u})\tilde{v}_{\mathrm{ice}}-\tilde{B}(\tilde{u})\binom{\tilde{h}}{\tilde{a}}-c_{\mathrm{cor}}\tilde{v}_{\mathrm{ice}}^{\perp}-g\nabla_{\mathrm{H}}H+\frac{1}{\rho_{\mathrm{ice}}\tilde{h}}\tau_{\mathrm{ice}}(\tilde{v}_{\mathrm{ice}}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{h}&=-\tilde{h}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}+S_{\mathrm{h}}(\tilde{h},\tilde{a}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{a}&=-\tilde{a}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}+S_{\mathrm{a}}(\tilde{h},\tilde{a}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \tilde{v}_{\mathrm{ice}}&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\Omega,\\ \tilde{v}_{\mathrm{ice}}(0)&=v_{\mathrm{ice},0},\enspace\tilde{h}(0)=h_{0},\enspace\tilde{a}(0)=a_{0},&&\enspace\text{in}\enspace\Omega.\end{aligned}\right.$
We now reformulate the main result in terms of the transformed system. The first task is then to verify this theorem, and in the end, we will deduce the assertion of Theorem 2.1 therefrom.
Theorem 3.1.
Let $p,q\in(1,\infty)$ satisfy (2.6), let $u_{0}\in V$, with $V$ as in (2.8), and recall the spaces $\mathrm{X}_{0}$ and $\mathrm{X}_{1}$ from (2.4). Additionally, suppose that the external terms $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}$, $H$ and $f_{\mathrm{gr}}$ fulfill Section 2. Then there exists $T>0$ such that (3.8) has a unique solution $\tilde{u}=(\tilde{v}_{\mathrm{ice}},\tilde{h},\tilde{a})$ with
$\tilde{u}\in\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1})\cap\mathrm{C}([0,T];V).$
Next, we study the continuous dependence of the solution on the initial data.
[…]
Under the assumptions of Theorem 3.1, let $u_{0}\in V$. Then there are $T>0$ and $r>0$ such that $\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r)\subset V$, and for all $u_{1}\in\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r)$, there exists a unique solution $\tilde{u}=\tilde{u}(\cdot,u_{1})$ to (3.8) on $(0,T)$ in the regularity class as in Theorem 3.1. Moreover, there exists a constant $C>0$ such that for all $u_{1}$, $u_{2}\in\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r)$, the respective solutions to (3.8) satisfy $\|\tilde{u}(\cdot,u_{1})-\tilde{u}(\cdot,u_{2})\|_{\mathbb{E}_{1}}\leq C\cdot\|u_{1}-u_{2}\|_{\mathrm{X}_{\gamma}}$.
We finish this section with a blow-up criterion for the solution $\tilde{u}$ in the Lagrangian formulation.
[…]
Under the assumptions of Theorem 3.1, the solution $\tilde{u}$ to (3.8) exists on a maximal time interval of existence $J(u_{0})=[0,t_{+}(u_{0}))$, where $t_{+}(u_{0})$ is characterized by
  1. (a)
    global existence, so $t_{+}(u_{0})=\infty$,
  2. (b)
    $\liminf_{t\to t_{+}(u_{0})}\operatorname{dist}_{\mathrm{X}_{\gamma}}(\tilde{u}(t),\partial V)=0$, or
  3. (c)
    $\lim_{t\to t_{+}(u_{0})}\tilde{u}(t)$ does not exist in $\mathrm{X}_{\gamma}$.
4. Linear theory
In this section, we address the linear theory. More precisely, we establish the boundedness of the $\mathcal{H}^{\infty}$-calculus of the operator matrix associated to the linearization of the parabolic-hyperbolic sea ice model. This result is of independent interest. As a corollary, we especially obtain the maximal $\mathrm{L}^{p}$-regularity for the fixed point argument.
First, we recall the $\mathrm{L}^{q}$-realization of the linearized Hibler operator as well as the bounded $\mathcal{H}^{\infty}$-calculus that it admits. For $u_{0}\in\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{C}^{0,\alpha}(\overline{\Omega})\times\mathrm{C}^{0,\alpha}(\overline{\Omega})$, with $\alpha>0$ and $h_{0}\geq\kappa$, we recall the linearized Hibler operator $\mathbb{A}^{\mathrm{H}}(u_{0})$ from (3.5). The $\mathrm{L}^{q}$-realization of the linearized Hibler operator subject to Dirichlet boundary conditions on $\partial\Omega$ is then given by
(4.1) $\left[A_{\mathrm{D}}^{\mathrm{H}}(u_{0})\right]v_{\mathrm{ice}}\coloneqq\left[\mathbb{A}^{\mathrm{H}}(u_{0})\right]v_{\mathrm{ice}},\enspace\text{with}\enspace\mathrm{D}(A_{\mathrm{D}}^{\mathrm{H}}(u_{0}))\coloneqq\mathrm{W}^{2,q}(\Omega)^{2}\cap\mathrm{W}_{0}^{1,q}(\Omega)^{2}.$
The following result can be obtained analogously as in [6, Section 4]. Let us observe that the reduced regularity in the $h$- and $a$-component does not affect the arguments as the coefficients remain Hölder-continuous, because they depend on $h$ and $a$ in a smooth way.
[…]
Let $q\in(1,\infty)$ as well as $u_{0}\in\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{C}^{0,\alpha}(\overline{\Omega})\times\mathrm{C}^{0,\alpha}(\overline{\Omega})$ for some $\alpha>0$ and with $h_{0}\geq\kappa$, and consider $A_{\mathrm{D}}^{\mathrm{H}}(u_{0})$ as defined in (4.1). Then there exists $\omega_{1}\in\mathbb{R}$ such that $-A_{\mathrm{D}}^{\mathrm{H}}(u_{0})+\omega\in\mathcal{H}^{\infty}(\mathrm{L}^{q}(\Omega)^{2})$ with $\mathcal{H}^{\infty}$-angle $\phi_{-A_{\mathrm{D}}^{\mathrm{H}}(u_{0})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$ for all $\omega>\omega_{1}$.
Note that the latter lemma especially implies the maximal $\mathrm{L}^{p}$-regularity of $-A_{\mathrm{D}}^{\mathrm{H}}(u_{0})+\omega$, see for example [11, Section 4.4]. In fact, mimicking the procedure in [6, Section 4], one can show the maximal regularity of the Hibler operator when merely assuming $u_{0}\in\mathrm{C}^{1}(\overline{\Omega})^{2}\times\mathrm{C}(\overline{\Omega})\times\mathrm{C}(\overline{\Omega})$.
Next, let $f_{1}\colon(0,T)\times\Omega\to\mathbb{R}^{2}$ and $f_{2}$, $f_{3}\colon(0,T)\times\Omega\to\mathbb{R}$. Besides, we consider suitable initial data $u_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})$ as well as $u_{1}\in\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{C}^{0,\alpha}(\overline{\Omega})\times\mathrm{C}^{0,\alpha}(\overline{\Omega})$, for $\alpha>0$, and with $h_{1}\geq\kappa$ and $a_{1}\in(0,1)$. For $\omega\geq 0$, the linear problem under investigation is given by
(4.2) $\partial_{t}\tilde{v}_{\mathrm{ice}}-(\mathbb{A}^{\mathrm{H}}(u_{1})-\omega)\tilde{v}_{\mathrm{ice}}+B_{1}(u_{1})\binom{\tilde{h}}{\tilde{a}}&=f_{1},&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{h}+h_{1}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}+\omega\tilde{h}&=f_{2},&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{a}+a_{1}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}+\omega\tilde{a}&=f_{3},&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \tilde{v}_{\mathrm{ice}}&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\Omega,\\ \tilde{v}_{\mathrm{ice}}(0)=v_{\mathrm{ice},0},\enspace\tilde{h}(0)=h_{0},\enspace\tilde{a}(0)&=a_{0},&&\enspace\text{in}\enspace\Omega,\end{aligned}\right.$
where $B_{1}$ is as defined by $B_{1}(u_{1})\binom{\tilde{h}}{\tilde{a}}=\frac{\partial_{h}P(h_{1},a_{1})}{2\rho_{\mathrm{ice}}h_{1}}\nabla_{\mathrm{H}}\tilde{h}+\frac{\partial_{a}P(h_{1},a_{1})}{2\rho_{\mathrm{ice}}h_{1}}\nabla_{\mathrm{H}}\tilde{a}$. The operator matrix corresponding to the linearized problem (4.2) for $\omega=0$ is given by
(4.3) $A(u_{1})\coloneqq\begin{pmatrix}-A_{\mathrm{D}}^{\mathrm{H}}(u_{1})&\frac{\partial_{h}P(h_{1},a_{1})}{2\rho_{\mathrm{ice}}h_{1}}\nabla_{\mathrm{H}}&\frac{\partial_{a}P(h_{1},a_{1})}{2\rho_{\mathrm{ice}}h_{1}}\nabla_{\mathrm{H}}\\ h_{1}\mathrm{div}_{\mathrm{H}}\,&0&0\\ a_{1}\mathrm{div}_{\mathrm{H}}\,&0&0\end{pmatrix}.$
The main result of this section asserts that, up to a shift, the operator matrix $A(u_{1})$ from (4.3) admits a bounded $\mathcal{H}^{\infty}$-calculus on $\mathrm{X}_{0}$. As a preparation, we briefly recall some theory on the $\mathcal{H}^{\infty}$-calculus for operator matrices. To this end, we first invoke the notion of a diagonally dominant block operator matrix.
[…]
Consider Banach spaces $\mathrm{X}$ and $\mathrm{Y}$ as well as linear operators $A\colon\mathrm{D}(A)\subset\mathrm{X}\to\mathrm{X}$, $B\colon\mathrm{D}(B)\subset\mathrm{Y}\to\mathrm{X}$, $C\colon\mathrm{D}(C)\subset\mathrm{X}\to\mathrm{Y}$ and $D\colon\mathrm{D}(D)\subset\mathrm{Y}\to\mathrm{Y}$. Moreover, set $\mathrm{Z}\coloneqq\mathrm{X}\times\mathrm{Y}$. If
  1. (a)
    the operators $A$ and $D$ are closed, linear and densely defined, and
  2. (b)
    it holds that $C$ is relatively $A$-bounded and $B$ is relatively $D$-bounded, i. e., $\mathrm{D}(D)\subset\mathrm{D}(B)$, $\mathrm{D}(A)\subset\mathrm{D}(C)$, and there are constants $c_{A}$, $c_{D}$, $C\geq 0$ with
    $\|Cx\|_{\mathrm{X}}$ $\leq c_{A}\cdot\|Ax\|_{\mathrm{X}}+L\cdot\|x\|_{\mathrm{X}}\enspace\text{for all}\enspace x\in\mathrm{D}(A),\enspace\text{and}\enspace$
    $\|By\|_{\mathrm{X}}$ $\leq c_{D}\cdot\|Dy\|_{\mathrm{Y}}+L\cdot\|y\|_{\mathrm{Y}}\enspace\text{for all}\enspace y\in\mathrm{D}(D),$
then the operator matrix $\mathcal{A}\colon\mathrm{D}(\mathcal{A})\coloneqq\mathrm{D}(A)\times\mathrm{D}(D)\subset\mathrm{Z}\to\mathrm{Z}$ defined by
(4.4) $\mathcal{A}\binom{x}{y}\coloneqq\begin{pmatrix}A&B\\ C&D\end{pmatrix}\binom{x}{y},\enspace\text{for}\enspace\binom{x}{y}\in\mathrm{D}(\mathcal{A}),$
is called diagonally dominant.
The following lemma tailored to our setting is a direct consequence of [1, Corollary 7.2].
[…]
Let $\mathcal{A}$ as defined in (4.4) be diagonally dominant in the sense of Section 4, and assume that $A$ is sectorial on $\mathrm{X}$ with spectral angle $\phi_{A}\in[0,\pi)$ as well as $D\in\mathcal{L}(\mathrm{Y})$. If in addition, $A\in\mathcal{H}^{\infty}(\mathrm{X})$ with $\phi_{A}^{\infty}\in[0,\pi)$, then for every $\phi\in(\phi_{A}^{\infty},\pi)$, there is $\omega\geq 0$ such that $\mathcal{A}+\omega\in\mathcal{H}^{\infty}(\mathrm{Z})$ and $\phi_{\mathcal{A}+\omega}^{\infty}\leq\phi$.
Let us note that the diagonal dominance of $\mathcal{A}$ together with the assumption $D\in\mathcal{L}(\mathrm{Y})$ already requires the boundedness of the operator $B$, i. e., $B\in\mathcal{L}(\mathrm{Y},\mathrm{X})$.
We are now in the position to address the main result of this section on the $\mathcal{H}^{\infty}$-calculus of $A(u_{1})$.
Theorem 4.1.
Let $q\in(2,\infty)$ and $u_{1}\in\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)$ for some $\alpha>0$ and with $h_{1}\geq\kappa$. Then there is $\omega_{0}\in\mathbb{R}$ such that for the operator matrix $A(u_{1})$ from (4.3), it holds that $A(u_{1})+\omega\in\mathcal{H}^{\infty}(\mathrm{X}_{0})$ with $\phi_{A(u_{1})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$ for all $\omega>\omega_{0}$.
Proof.
The idea of the proof is to make use of Section 4. At this stage, the choice of the anisotropic ground space $\mathrm{X}_{0}$ as introduced in (2.4) is crucial in order to establish the diagonal dominance of $A(u_{1})$.
For this purpose, we recall from Section 4 that for $\omega>\omega_{1}$, we have $-A_{\mathrm{D}}^{\mathrm{H}}(u_{1})+\omega\in\mathcal{H}^{\infty}(\mathrm{L}^{q}(\Omega)^{2})$ with $\phi_{-A_{\mathrm{D}}^{\mathrm{H}}(u_{1})+\omega}^{\infty}<\nicefrac{{\pi}}{{2}}$ upon observing that $\mathrm{W}^{1,q}(\Omega)\hookrightarrow\mathrm{C}^{0,\beta}(\overline{\Omega})$ for some $\beta>0$ thanks to $q>2$. On the other hand, for such $\omega$, the operator $\operatorname{diag}(\omega\operatorname{Id},\omega\operatorname{Id})$ is clearly bounded on $\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)$. It thus remains to handle the off-diagonal terms. With regard to the remark following Section 4, we have to show the boundedness of the terms above the diagonal. The shape of $P(h_{1},a_{1})$ from (2.1) as well as the assumptions on $u_{1}$ directly yield the existence of $C>0$ such that
$\left\|\frac{\partial_{h}P(h_{1},a_{1})}{2\rho_{\mathrm{ice}}h_{1}}\nabla_{\mathrm{H}}h\right\|_{\mathrm{L}^{q}(\Omega)}\leq C\cdot\|h\|_{\mathrm{W}^{1,q}(\Omega)},\enspace\text{for}\enspace h\in\mathrm{W}^{1,q}(\Omega),$
and likewise for the corresponding $a$-term. Next, let $v_{\mathrm{ice}}\in\mathrm{D}(A_{\mathrm{D}}^{\mathrm{H}}(u_{1}))=\mathrm{W}^{2,q}(\Omega)^{2}\cap\mathrm{W}_{0}^{1,q}(\Omega)$. Employing the Banach algebra structure of $\mathrm{W}^{1,q}(\Omega)$ by $q>2$, and using the equivalence of the norm of $\mathrm{W}^{2,q}(\Omega)^{2}$ with the graph norm of $-A_{\mathrm{D}}^{\mathrm{H}}(u_{1})+\omega$ for $\omega>0$ sufficiently large, we find that
$\|h_{1}\mathrm{div}_{\mathrm{H}}\,v_{\mathrm{ice}}\|_{\mathrm{W}^{1,q}(\Omega)}\leq C\cdot\|v_{\mathrm{ice}}\|_{\mathrm{W}^{2,q}(\Omega)}\leq C\cdot\|(-A_{\mathrm{D}}^{\mathrm{H}}(u_{1})+\omega)v_{\mathrm{ice}}\|_{\mathrm{L}^{q}(\Omega)}.$
The other term below the diagonal can be handled analogously. In total, we infer that the operator matrix $A(u_{1})+\omega$, with $A(u_{1})$ from (4.3), is diagonally dominant. Thus, for $\omega_{0}\geq\omega_{1}$, we deduce the assertion of the theorem from Section 4. ∎
Having proved the $\mathcal{H}^{\infty}$-calculus of the operator matrix associated to the linearization, we now focus on the maximal $\mathrm{L}^{p}$-regularity of the linearized problem (4.2). To this end, let us briefly introduce the so-called data space $\mathbb{E}_{0}$ and the maximal regularity space $\mathbb{E}_{1}$. With $\mathrm{X}_{0}$ and $\mathrm{X}_{1}$ from (2.4), we set
(4.5) $\mathbb{E}_{0}\coloneqq\mathrm{L}^{p}(0,T;\mathrm{X}_{0})\enspace\text{and}\enspace\mathbb{E}_{1}\coloneqq\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1}).$
The maximal regularity result below is a consequence of Theorem 4.1, since a bounded $\mathcal{H}^{\infty}$-calculus with angle smaller than $\nicefrac{{\pi}}{{2}}$ implies the maximal $\mathrm{L}^{p}$-regularity on UMD spaces, see e. g. [11, Section 4], and the present ground space $\mathrm{X}_{0}$ from (2.4) is in particular a UMD space.
[…]
Let $T\in(0,\infty]$, $p\in(1,\infty)$, $q\in(2,\infty)$ with $\nicefrac{{2}}{{p}}+\nicefrac{{1}}{{q}}\neq 2$, and for $u_{1}=(v_{\mathrm{ice},1},h_{1},a_{1})$, assume $u_{1}\in\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)$, with $\alpha>0$, $h_{1}\geq\kappa$ and $a_{1}\in(0,1)$. Furthermore, for the time trace space $\mathrm{X}_{\gamma}$ and the data space $\mathbb{E}_{0}$ as in (2.5) and (4.5), consider $u_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in\mathrm{X}_{\gamma}$ and $(f_{1},f_{2},f_{3})\in\mathbb{E}_{0}$. Then there exists $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, the linearized problem (4.2) has a unique solution $\tilde{u}=(\tilde{v}_{\mathrm{ice}},\tilde{h},\tilde{a})\in\mathbb{E}_{1}$, with $\mathbb{E}_{1}$ as introduced in (4.5). Besides, there is a constant $C_{\mathrm{MR}}>0$ so that the unique solution $\tilde{u}$ of (4.2) satisfies
$\|\tilde{u}\|_{\mathbb{E}_{1}}\leq C_{\mathrm{MR}}\left(\|(f_{1},f_{2},f_{3})\|_{\mathbb{E}_{0}}+\|u_{0}\|_{\mathrm{X}_{\gamma}}\right).$
[…]
The assertion of Section 4 can also be obtained by a direct perturbation argument when only assuming $u_{1}\in\mathrm{C}^{1}(\overline{\Omega})^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)$, see also [5, Proposition 6.2.1].
In the sequel, we fix $\omega>\omega_{0}$ for $\omega_{0}>0$ from Section 4. The maximal regularity constant $C_{\mathrm{MR}}>0$ can in particular be chosen independent of $T>0$ provided homogeneous initial values are considered. In order reduce the considerations to this situation, we now discuss the so-called reference solution capturing the initial values. More precisely, for $u_{0}\in V$, the resulting system reads as
(4.6) $\partial_{t}\tilde{v}_{\mathrm{ice}}-(\mathbb{A}^{\mathrm{H}}(u_{0})-\omega)\tilde{v}_{\mathrm{ice}}+B_{1}(u_{0})\binom{\tilde{h}}{\tilde{a}}&=0,&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{h}+h_{0}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}+\omega\tilde{h}&=0,&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\tilde{a}+a_{0}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}+\omega\tilde{a}&=0,&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \tilde{v}_{\mathrm{ice}}&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\Omega,\\ \tilde{v}_{\mathrm{ice}}(0)=v_{\mathrm{ice},0},\enspace\tilde{h}(0)=h_{0},\enspace\tilde{a}(0)&=a_{0},&&\enspace\text{in}\enspace\Omega.\end{aligned}\right.$
The result below on the existence of a unique reference solution to (4.6) follows from Section 4.
[…]
Let $p,q\in(1,\infty)$ satisfy (2.6), and consider $0<T\leq\infty$ and $u_{0}=(v_{\mathrm{ice},0},h_{0},a_{0})\in V$, with $V$ from (2.8). Then (4.6) admits a unique solution $u_{0}^{*}=(v_{\mathrm{ice},0}^{*},h_{0}^{*},a_{0}^{*})\in\mathbb{E}_{1}$, with $\mathbb{E}_{1}$ from (4.5).
The importance of the reference solution $u_{0}^{*}$ for the arguments later on calls for additional remarks.
[…]
  1. (a)
    The reference solution $u_{0}^{*}$ results from an application of the semigroup generated by $A(u_{0})+\omega$, with $A(u_{0})$ as defined in (4.3), to the initial values $u_{0}$, so $u_{0}^{*}(t)=\mathrm{e}^{-(A(u_{0})+\omega)(t)}u_{0}$.
  2. (b)
    The norm of $u_{0}^{*}$ will be denoted by $C_{T}^{*}\coloneqq\|u_{0}^{*}\|_{\mathbb{E}_{1}}$, and we observe that $C_{T}^{*}\to 0$ as $T\to 0$.
  3. (c)
    From (a), it follows that $\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$ as $T\to 0$. In particular, as $V\subset\mathrm{X}_{\gamma}$ is open, there is a sufficiently small $r_{0}>0$ with $\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r_{0})\subset V$ for given $u_{0}\in V$. Thus, $u_{0}^{*}(t)\in V$ holds for all $t\in[0,T_{0}]$, for $T_{0}>0$ sufficiently small, since
    (4.7) $\sup_{t\in[0,T_{0}]}\|u_{0}^{*}(t)-u_{0}\|_{\mathrm{X}_{\gamma}}\leq\frac{r_{0}}{2}.$
5. Proof of the main result
This final section presents the proof of the main result. For this, we first elaborate on the fixed point argument, then provide the required estimates and conclude the proof of Theorem 3.1 by the contraction mapping principle. The final step is to deduce the proof of Theorem 2.1 from there.
5.1. Reformulation as a fixed point problem
The aim of this subsection is to reformulate the task of finding a solution to the transformed system (3.8) as a fixed point problem. Thus, consider a solution $\tilde{u}=(\tilde{v}_{\mathrm{ice}},\tilde{h},\tilde{a})$ to (3.8), recall the reference solution $u_{0}^{*}=(v_{\mathrm{ice},0}^{*},h_{0}^{*},a_{0}^{*})$ from Section 4 and define $\widehat{u}=(\widehat{v}_{\mathrm{ice}},\widehat{h},\widehat{a})$ by $\widehat{u}\coloneqq\tilde{u}-u_{0}^{*}$. Then $\widehat{u}$ solves
(5.1) $\partial_{t}\widehat{v}_{\mathrm{ice}}-(\mathbb{A}^{\mathrm{H}}(u_{0})-\omega)\widehat{v}_{\mathrm{ice}}+B_{1}(u_{0})\binom{\widehat{h}}{\widehat{a}}&=F_{1}(\widehat{u}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\widehat{h}+h_{0}\mathrm{div}_{\mathrm{H}}\,\widehat{v}_{\mathrm{ice}}+\omega\widehat{h}&=F_{2}(\widehat{u}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \partial_{t}\widehat{a}+a_{0}\mathrm{div}_{\mathrm{H}}\,\widehat{v}_{\mathrm{ice}}+\omega\widehat{a}&=F_{3}(\widehat{u}),&&\enspace\text{in}\enspace(0,T)\times\Omega,\\ \widehat{v}_{\mathrm{ice}}&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\Omega,\\ \widehat{v}_{\mathrm{ice}}(0)=0,\enspace\widehat{h}(0)=0,\enspace\widehat{a}(0)&=0,&&\enspace\text{in}\enspace\Omega,\end{aligned}\right.$
for the right-hand sides
(5.2) $F_{1}(\widehat{u})$ $\coloneqq\bigl{(}\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}\tilde{v}_{\mathrm{ice}}-\bigl{(}\tilde{B}(\tilde{u})-B_{1}(u_{0})\bigr{)}\binom{\tilde{h}}{\tilde{a}}+\omega\tilde{v}_{\mathrm{ice}}-c_{\mathrm{cor}}\tilde{v}_{\mathrm{ice}}^{\perp}-g\nabla_{\mathrm{H}}H$
$\quad+\frac{1}{\rho_{\mathrm{ice}}\tilde{h}}\bigl{(}\tau_{\mathrm{atm}}+\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice}})\bigr{)},$
$F_{2}(\widehat{u})$ $\coloneqq h_{0}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}-\tilde{h}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}+\omega\tilde{h}+S_{\mathrm{h}}(\tilde{h},\tilde{a})\enspace\text{and}\enspace$
$F_{3}(\widehat{u})$ $\coloneqq a_{0}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}-\tilde{a}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}+\omega\tilde{a}+S_{\mathrm{a}}(\tilde{h},\tilde{a}).$
In the following, we denote by $\prescript{}{0}{\mathbb{E}_{1}}$ the elements in the maximal regularity space $\mathbb{E}_{1}$ from (4.5) with homogeneous initial values, so $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ fulfills $\widehat{u}(0)=0$. With regard to (5.1), the solution to the transformed problem $\tilde{u}$ results from the addition of $\widehat{u}\in\mathbb{E}_{1}$ and the reference solution $u_{0}^{*}$. We demand that this solution also only takes values in the physically relevant range, so it should hold that $\tilde{u}(t)\in V$ for all $t\in[0,T_{1}]$, where $T_{1}>0$ will be determined in the sequel. Recall that given $u_{0}\in V$, there exists $r_{0}>0$ so that $\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r_{0})\subset V$. For $T_{0}>0$ as in Section 4, we especially have $u_{0}^{*}(t)\in V$ for all $t\in[0,T_{0}]$. Moreover, we deduce from [2, Theorem III.4.10.2] the embedding
(5.3) $\prescript{}{0}{\mathbb{E}_{1}}\hookrightarrow\mathrm{BUC}([0,T_{0}];\mathrm{X}_{\gamma}).$
In the above, $\mathrm{BUC}([0,T_{0}];\mathrm{X}_{\gamma})$ represents the bounded and uniformly continuous functions in time with values in $\mathrm{X}_{\gamma}$. Combining (4.7) and the choice $R_{0}\leq\nicefrac{{r_{0}}}{{2C}}$ for the embedding constant $C>0$ from (5.3), we conclude for $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R_{0}$ that
$\sup_{t\in[0,T_{0}]}\|\tilde{u}(t)-u_{0}\|_{\mathrm{X}_{\gamma}}\leq C\cdot\|\widehat{u}\|_{\mathbb{E}_{1}}+\sup_{t\in[0,T_{0}]}\|u_{0}^{*}(t)-u_{0}\|_{\mathrm{X}_{\gamma}}\leq\frac{r_{0}}{2}+\frac{r_{0}}{2}\leq r_{0}.$
The preceding discussion can be summarized in the lemma below.
[…]
Let $u_{0}\in V$, $T_{0}>0$ as introduced in Section 4 and $0<R_{0}\leq\nicefrac{{r_{0}}}{{2C}}$, with $C>0$ representing the embedding constant from (5.3), and $r_{0}>0$ is such that $\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r_{0})\subset V$. Consider $T\in(0,T_{0}]$, $R\in(0,R_{0}]$ and $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$, where $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$, and $u_{0}^{*}\in\mathbb{E}_{1}$ denotes the reference solution from Section 4. Then $\tilde{u}(t)\in V$ for all $t\in[0,T]$.
Next, we describe the fixed point argument in more details. Let $p,q\in(1,\infty)$ be such that (2.6) is valid, and consider $T_{0}>0$ and $R_{0}>0$ from Subsection 5.1. For $T\in(0,T_{0}]$ and $R\in(0,R_{0}]$, we then define
(5.4) $\mathcal{K}_{T}^{R}\coloneqq\{\overline{u}\in\prescript{}{0}{\mathbb{E}_{1}}:\|\overline{u}\|_{\mathbb{E}_{1}}\leq R\}\enspace\text{and}\enspace\Phi_{T}^{R}\colon\mathcal{K}_{T}^{R}\to\prescript{}{0}{\mathbb{E}_{1}},\enspace\text{with}\enspace\Phi_{T}^{R}(\overline{u})\coloneqq\widehat{u},$
where $\widehat{u}$ represents the unique solution to (5.1) with right-hand sides $F_{1}(\overline{u})$, $F_{2}(\overline{u})$ and $F_{3}(\overline{u})$ as introduced in (5.2). By the maximal regularity result, Section 4, the map $\Phi_{T}^{R}$ is well-defined provided the terms on the right-hand side lie in the data space $\mathbb{E}_{0}$ as introduced in (4.5).
5.2. Estimates of the nonlinear terms
This subsection presents the estimates of the nonlinear terms $F_{1}$, $F_{2}$ and $F_{3}$. For this, we require some further preparation, and we start by discussing properties of the transformation to Lagrangian coordinates. Recall that
(5.5) $X(t,y_{\mathrm{H}})=y_{\mathrm{H}}+\int_{0}^{t}\tilde{v}_{\mathrm{ice}}(s,y_{\mathrm{H}})\,\mathrm{d}s.$
Moreover, for the inverse $Y(t,\cdot)$ of $X(t,\cdot)$, we have $\nabla_{\mathrm{H}}Y(t,X(t,y_{\mathrm{H}}))=[\nabla_{\mathrm{H}}X]^{-1}(t,y_{\mathrm{H}})$. Aiming to track the dependence of the diffeomorphisms $X$ and $Y$ on $\tilde{v}_{\mathrm{ice}}$, we also use the notation $X_{\tilde{v}_{\mathrm{ice}}}$ and $Y_{\tilde{v}_{\mathrm{ice}}}$. In the sequel, $\tilde{v}_{\mathrm{ice}}$ results from the addition of $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ and the reference solution $u_{0}^{*}$ from Section 4. We also introduce the notation $\mathbb{E}_{1}=\mathbb{E}_{1}^{v}\times\mathbb{E}_{1}^{h}\times\mathbb{E}_{1}^{a}$, and we conclude that
(5.6) $\|\tilde{u}\|_{\mathbb{E}_{1}}\leq R+C_{T}^{*}\leq R_{0}+C_{T_{0}}^{*},\enspace\text{so}\enspace\|\tilde{v}_{\mathrm{ice}}\|_{\mathbb{E}_{1}^{v}}\leq R_{0}+C_{T_{0}}^{*}.$
Given $X=X_{\tilde{v}_{\mathrm{ice}}}$ as in (5.5) and $p^{\prime}\in(1,\infty)$ representing the Hölder conjugate of $p$, we deduce from Hölder’s inequality and (5.6) that
(5.7) $\sup_{t\in[0,T]}\|\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}-\operatorname{Id}_{2}\|_{\mathrm{W}^{1,q}(\Omega)}\leq C\int_{0}^{T}\|\nabla_{\mathrm{H}}\tilde{v}_{\mathrm{ice}}(t,\cdot)\|_{\mathrm{W}^{1,q}(\Omega)}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}(R_{0}+C_{T_{0}}^{*}).$
By virtue of the embedding $\mathrm{W}^{1,q}(\Omega)\hookrightarrow\mathrm{L}^{\infty}(\Omega)$, thanks to $q>2$, we infer that there is $T_{0}^{\prime}>0$ small with
(5.8) $\sup_{t\in[0,T_{0}^{\prime}]}\|\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}-\operatorname{Id}_{2}\|_{\mathrm{L}^{\infty}(\Omega)}\leq\frac{1}{2}.$
As a result, $\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}(t,\cdot)$ is invertible for all $t\in[0,T_{0}^{\prime}]$, and $\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}(t,\cdot)$ thus exists on $[0,T_{0}^{\prime}]$. It also follows from $\partial_{t}\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}(t,y_{\mathrm{H}})=\nabla_{\mathrm{H}}\tilde{v}_{\mathrm{ice}}(t,y_{\mathrm{H}})$ and (5.6) that
$\|\partial_{t}\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}(t,\cdot)\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq C\cdot\|\tilde{v}_{\mathrm{ice}}\|_{\mathbb{E}_{1}^{v}}\leq C(R_{0}+C_{T_{0}}^{*}).$
Similarly, for $\tilde{u}_{i}=\widehat{u}_{i}+u_{0}^{*}$, with $\widehat{u}_{i}\in\prescript{}{0}{\mathbb{E}_{1}}$ and $u_{0}^{*}$ representing the reference solution from Section 4, we denote the diffeomorphisms related to $\tilde{v}_{\mathrm{ice},i}$ by $X^{(i)}$ and $Y^{(i)}$ for $i=1,2$. As in (5.7), it follows that
$\sup_{t\in[0,T]}\|\nabla_{\mathrm{H}}X^{(1)}-\nabla_{\mathrm{H}}X^{(2)}\|_{\mathrm{L}^{\infty}(\Omega)}+\sup_{t\in[0,T]}\|\nabla_{\mathrm{H}}X^{(1)}-\nabla_{\mathrm{H}}X^{(2)}\|_{\mathrm{W}^{1,q}(\Omega)}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}.$
In addition, the identity $\partial_{t}(\nabla_{\mathrm{H}}X^{(1)}-\nabla_{\mathrm{H}}X^{(2)})=\nabla_{\mathrm{H}}(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})$ leads to
$\|\partial_{t}\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq C\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}.$
We collect these considerations in the lemma below, where we additionally include estimates of $\nabla_{\mathrm{H}}Y$.
[…]
Let $p,q\in(1,\infty)$ satisfy (2.6), and consider $T\in(0,T_{1}]$, where $T_{1}\coloneqq\min\{T_{0},T_{0}^{\prime}\}>0$, for $T_{0}>0$ chosen in (4.7) and $T_{0}^{\prime}>0$ related to (5.8). Moreover, let $R\in(0,R_{0}]$, with $R_{0}>0$ as in Subsection 5.1, and recall the reference solution $u_{0}^{*}$ from Section 4.
  1. (a)
    Consider $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$, where $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$. Then for $X_{\tilde{v}_{\mathrm{ice}}}$ as in (5.5) as well as $\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}\coloneqq[\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}]^{-1}$, whose existence is guaranteed by (5.8), for some $C>0$, we get
    $\|\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\enspace\text{and}\enspace$
    $\|\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C.$
  2. (b)
    Let $\tilde{u}_{i}=\widehat{u}_{i}+u_{0}^{*}$, with $\widehat{u}_{i}\in\prescript{}{0}{\mathbb{E}_{1}}$ and $\|\widehat{u}_{i}\|_{\mathbb{E}_{1}}\leq R$, $i=1,2$. Then for $X^{(1)}$, $X^{(2)}$, $Y^{(1)}$ and $Y^{(2)}$, there is a constant $C>0$ such that
    $\|\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)})\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)})\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}\enspace\text{and}\enspace$
    $\|\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}.$
Proof.
The estimates of $\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}$ follow from the above considerations. Thanks to (2.6), the spaces $\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))$ and $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))$ have the Banach algebra structure, so the norm of products can be estimated by the product of the norms. From the estimates of $\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}$, it then follows that
$\|\det\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\det\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\enspace\text{and}\enspace$
$\|\operatorname{Cof}\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{W}^{1,p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|\operatorname{Cof}\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C.$
With regard to (5.8), we have $\det\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}\geq C>0$ on $(0,T)\times\Omega$ if $T\leq T_{0}^{\prime}$, so the representation
$\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}=[\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}]^{-1}=\frac{1}{\det\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}}}(\operatorname{Cof}\nabla_{\mathrm{H}}X_{\tilde{v}_{\mathrm{ice}}})^{\top}$
leads to the desired estimates of $\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}$, showing (a). Concerning (b), we also remark that the estimates of $\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)})$ follow from the arguments given before this lemma, while for the estimates of $\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})$, we can proceed in a similar way as for (a). In fact, we use the relation
$\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})=\nabla_{\mathrm{H}}Y^{(1)}-\nabla_{\mathrm{H}}Y^{(2)}=-\nabla_{\mathrm{H}}Y^{(1)}(\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)}))\nabla_{\mathrm{H}}Y^{(2)}.$
Finally, the estimates follow by a concatenation of (a) and the estimates of $\nabla_{\mathrm{H}}(X^{(1)}-X^{(2)})$ in (b). ∎
Before estimating the nonlinear terms, we collect some further useful embeddings and relations. The assertion of (a) follows by [2, Theorem III.4.10.2], where the $T$-independence results from the existence of an extension operator with $T$-independent norm thanks to the homogeneous initial values. Aspect (b) is a direct consequence of Sobolev embeddings and (2.7), while (c) is implied by Hölder’s inequality.
[…]
Let $p,q\in(1,\infty)$ be such that (2.6) is valid, consider $T\in(0,T_{1}]$, where $T_{1}>0$ is as made precise in Subsection 5.2, and recall the trace space $\mathrm{X}_{\gamma}=\mathrm{X}_{\gamma}^{v}\times\mathrm{X}_{\gamma}^{h}\times\mathrm{X}_{\gamma}^{a}$ from (2.5).
  1. (a)
    We have $\prescript{}{0}{\mathbb{E}_{1}}\hookrightarrow\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})$, with $T$-independent embedding constant.
  2. (b)
    It holds that $\mathrm{BUC}([0,T];\mathrm{X}_{\gamma}^{v})\hookrightarrow\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2q}(\Omega)^{2})\hookrightarrow\mathrm{L}^{2p}(0,T;\mathrm{L}^{2q}(\Omega)^{2})$. In particular, for $T$-independent embedding constant, we obtain $\prescript{}{0}{\mathbb{E}_{1}^{v}}\hookrightarrow\mathrm{BUC}([0,T];\mathrm{X}_{\gamma}^{v})\hookrightarrow\mathrm{L}^{2p}(0,T;\mathrm{L}^{2q}(\Omega)^{2})$.
  3. (c)
    For $f\in\mathrm{L}^{\infty}(0,T)$, we get $\|f\|_{\mathrm{L}^{p}(0,T)}\leq T^{\nicefrac{{1}}{{p}}}\cdot\|f\|_{\mathrm{L}^{\infty}(0,T)}$, while for $f\in\mathrm{W}^{1,p}(0,T)$ with $f(0)=0$, we get $\|f\|_{\mathrm{L}^{\infty}(0,T)}\leq T^{\nicefrac{{1}}{{p^{\prime}}}}\cdot\|f\|_{\mathrm{W}^{1,p}(0,T)}$, and $p^{\prime}\in(1,\infty)$ denotes again the Hölder conjugate of $p$.
A combination of Subsection 5.2 and Subsection 5.2(c) leads to the following estimate of the difference of $\nabla_{\mathrm{H}}Y$ from $\operatorname{Id}_{2}$ as well as the difference $\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})$.
[…]
Let $p,q\in(1,\infty)$ fulfill (2.6), let $T\in(0,T_{1}]$ and $R\in(0,R_{0}]$, with $T_{1}>0$ from Subsection 5.2 and $R_{0}>0$ as in Subsection 5.1. Moreover, recall the reference solution $u_{0}^{*}$ from Section 4.
  1. (a)
    Let $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$ for $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$. Then for $X_{\tilde{v}_{\mathrm{ice}}}$ from (5.5), $\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}=[\nabla_{\mathrm{H}}X_{v_{\mathrm{ice}}}]^{-1}$ and the Hölder conjugate $p^{\prime}\in(1,\infty)$ of $p$, we have $\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y_{\tilde{v}_{\mathrm{ice}}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}$.
  2. (b)
    For $i=1,2$, consider $\tilde{u}_{i}=\widehat{u}_{i}+u_{0}^{*}$, where $\widehat{u}_{i}\in\prescript{}{0}{\mathbb{E}_{1}}$ and $\|\widehat{u}_{i}\|_{\mathbb{E}_{1}}\leq R$. Then for a constant $C>0$ and the Hölder conjugate $p^{\prime}\in(1,\infty)$ of $p$, the maps $Y^{(1)}$ and $Y^{(2)}$ satisfy the estimate $\|\nabla_{\mathrm{H}}(Y^{(1)}-Y^{(2)})\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}$.
Having the analytical tools at hand, we now start with the actual estimates of the nonlinear terms. In the sequel, we fix $T_{1}>0$ from Subsection 5.2 and $R_{0}>0$ from Subsection 5.1, and we consider $T\in(0,T_{1}]$ and $R\in(0,R_{0}]$. We begin with the quasilinear terms from the non-transformed equation.
[…]
Let $p,q\in(1,\infty)$ satisfy (2.6), and consider $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$ for $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$ and $u_{0}^{*}$ denoting the reference solution from Section 4. Then there exists a constant $C>0$ with
$\biggl{\|}\bigl{(}\mathbb{A}^{\mathrm{H}}(u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}\tilde{v}_{\mathrm{ice}}-\bigl{(}B_{1}(u_{0}^{*})-B_{1}(u_{0})\bigr{)}\binom{\tilde{h}}{\tilde{a}}\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq C\cdot\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}(R+C_{T}^{*})$
and
$\biggl{\|}\bigl{(}\mathbb{A}^{\mathrm{H}}(\tilde{u})-\mathbb{A}^{\mathrm{H}}(u_{0}^{*})\bigr{)}(\tilde{v}_{\mathrm{ice}})-\bigl{(}B_{1}(\tilde{u})-B_{1}(u_{0}^{*})\bigr{)}\binom{\tilde{h}}{\tilde{a}}\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CR(R+C_{T}^{*}).$
Proof.
Let $u_{1},u_{2}\in V$ and $w=(v_{\mathrm{ice}},h,a)\in\mathrm{X}_{1}$. Similarly as in [6, Section 6], additionally observing that the reduced regularity in the $h$- and $a$-component does not affect the arguments thanks to the smooth dependence of the coefficients on $\varepsilon$, $h$ and $a$, we deduce the existence of a constant $C_{A}>0$ such that
(5.9) $\biggl{\|}\bigl{(}\mathbb{A}^{\mathrm{H}}(u_{1})-\mathbb{A}^{\mathrm{H}}(u_{2})\bigr{)}v_{\mathrm{ice}}-\bigl{(}B_{1}(u_{1})-B_{1}(u_{2})\bigr{)}\binom{h}{a}\biggr{\|}_{\mathrm{L}^{q}(\Omega)}\leq C_{A}\cdot\|u_{1}-u_{2}\|_{\mathrm{X}_{\gamma}}\cdot\|w\|_{\mathrm{X}_{1}}.$
Recalling from Section 4(c) that $u_{0}^{*}(t)\in V$ for all $t\in[0,T]$ and $\tilde{u}(t)=\widehat{u}(t)+u_{0}^{*}(t)\in V$ for all $t\in[0,T]$, we can use (5.9) in the case that $u_{1}=u_{0}^{*}$, $u_{2}=u_{0}$ and $w=u_{0}^{*}$ to derive
$\quad\biggl{\|}\bigl{(}\mathbb{A}^{\mathrm{H}}(u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}v_{\mathrm{ice},0}^{*}-\bigl{(}B_{1}(u_{0}^{*})-B_{1}(u_{0})\bigr{)}\binom{h_{0}^{*}}{a_{0}^{*}}\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}$
$=\biggl{(}\int_{0}^{T}\biggl{\|}\bigl{(}\mathbb{A}^{\mathrm{H}}(u_{0}^{*}(t))-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}v_{\mathrm{ice},0}^{*}(t)-\bigl{(}B_{1}(u_{0}^{*}(t))-B_{1}(u_{0})\bigr{)}\binom{h_{0}^{*}(t)}{a_{0}^{*}(t)}\biggr{\|}_{\mathrm{L}^{q}(\Omega)}^{p}\,\mathrm{d}t\biggr{)}^{\nicefrac{{1}}{{p}}}$
$\leq C_{A}\biggl{(}\int_{0}^{T}\|u_{0}^{*}(t)-u_{0}\|_{\mathrm{X}_{\gamma}}^{p}\cdot\|u_{0}^{*}(t)\|_{\mathrm{X}_{1}}^{p}\,\mathrm{d}t\biggr{)}^{\nicefrac{{1}}{{p}}}$
$\leq C\cdot\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\cdot\|u_{0}^{*}\|_{\mathbb{E}_{1}}\leq C\cdot\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}C_{T}^{*}.$
This proves the first part of the assertion. The second part can be shown likewise. ∎
The following proposition establishes an estimate of the term $F_{1}$.
[…]
Let $p,q\in(1,\infty)$ satisfy (2.6) and $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$, with $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$, $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$ and $u_{0}^{*}$ representing the reference solution from Section 4. Moreover, suppose that $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}$, $\nabla_{\mathrm{H}}H$ and $f_{\mathrm{gr}}$ fulfill Section 2. Recall the $T$-independent maximal regularity constant $C_{\mathrm{MR}}>0$ from Section 4. Then there is $C_{F_{1}}(R,T)>0$, with $C_{F_{1}}(R,T)<\nicefrac{{R}}{{6C_{\mathrm{MR}}}}$ for $R>0$ and $T>0$ sufficiently small, such that $\|F_{1}(\widehat{u})\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq C_{F_{1}}(R,T)$.
Proof.
Before starting with the estimates, let us first comment on some commonly used embeddings and notation. We heavily rely on $\mathrm{W}^{1,p}(0,T)\hookrightarrow\mathrm{L}^{\infty}(0,T)$. In addition, we frequently use $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$ and often invoke the abbreviation $C_{T}^{*}$ of the norm of the reference solution. By $C_{0}\coloneqq\|u_{0}\|_{\mathrm{X}_{\gamma}}$, we denote the norm of the initial values in the sequel. Instead of $Y_{\tilde{v}_{\mathrm{ice}}}$, we write $Y$ throughout this proof, since we only consider $\tilde{u}=\widehat{u}+u_{0}^{*}$. First, Subsection 5.2(c) implies
$\left\|\omega\tilde{v}_{\mathrm{ice}}-c_{\mathrm{cor}}\tilde{v}_{\mathrm{ice}}^{\perp}\right\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq C(T^{\nicefrac{{1}}{{p}}}\cdot\|\widehat{v}_{\mathrm{ice}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{q}(\Omega))}+\|v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))})\leq C(T^{\nicefrac{{1}}{{p}}}R+C_{T}^{*}).$
By Section 2(b), we further get $\|-g\nabla_{\mathrm{H}}H\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq T^{\nicefrac{{1}}{{p}}}\cdot\|-g\nabla_{\mathrm{H}}H\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p}}}$.
For the subsequent considerations, it is useful to estimate the factor $\nicefrac{{1}}{{\rho_{\mathrm{ice}}(\widehat{h}+h_{0}^{*})}}$ in $\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))$. By virtue of Subsection 5.1, we especially obtain $\tilde{h}(t)=\widehat{h}(t)+h_{0}^{*}(t)>\kappa$ for all $t\in[0,T]$, yielding
(5.10) $\biggl{\|}\frac{1}{\rho_{\mathrm{ice}}\tilde{h}}\biggr{\|}_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\leq C.$
From (5.10), Subsection 5.2(c) and Section 2(a), we deduce that
$\biggl{\|}\frac{1}{\rho_{\mathrm{ice}}\tilde{h}}\tau_{\mathrm{atm}}\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p}}}\cdot\|V_{\mathrm{atm}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2q}(\Omega))}^{2}\leq CT^{\nicefrac{{1}}{{p}}}.$
Next, let us recall the well known estimate
(5.11) $\|u_{0}^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\leq C(\|u_{0}\|_{\mathrm{X}_{\gamma}}+\|u_{0}^{*}\|_{\mathbb{E}_{1}})\leq C(C_{0}+C_{T}^{*}),$
where the constant $C>0$ appearing in the estimate is independent of $T>0$. It then follows from (5.10), Subsection 5.2(c), Subsection 5.2(b), (5.11) and Section 2(a) that
$\biggl{\|}\frac{1}{\rho_{\mathrm{ice}}\tilde{h}}\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice}})\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}$ $\leq CT^{\nicefrac{{1}}{{p}}}\bigl{(}\|V_{\mathrm{ocn}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2q}(\Omega))}^{2}+\|\widehat{v}_{\mathrm{ice}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2q}(\Omega))}^{2}+\|v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{2q}(\Omega))}^{2}\bigr{)}$
$\leq CT^{\nicefrac{{1}}{{p}}}\left(1+R^{2}+C_{0}^{2}+(C_{T}^{*})^{2}\right).$
The term to be still estimated is $\bigl{(}\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}\tilde{v}_{\mathrm{ice}}-\bigl{(}\tilde{B}(\tilde{u})-B_{1}(u_{0})\bigr{)}\binom{\tilde{h}}{\tilde{a}}$. This comes down to
(5.12) $\bigl{(}\tilde{\mathbb{A}}^{\mathrm{H}}(\widehat{u}+u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(\widehat{u}+u_{0}^{*})\bigr{)}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})-\bigl{(}\tilde{B}(\widehat{u}+u_{0}^{*})-B_{1}(\widehat{u}+u_{0}^{*})\bigr{)}\binom{\widehat{h}+h_{0}^{*}}{\widehat{a}+a_{0}^{*}}\enspace\text{and}\enspace$
$\bigl{(}\mathbb{A}^{\mathrm{H}}(\widehat{u}+u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(u_{0}^{*})\bigr{)}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})-\bigl{(}B_{1}(\widehat{u}+u_{0}^{*})-B_{1}(u_{0}^{*})\bigr{)}\binom{\widehat{h}+h_{0}^{*}}{\widehat{a}+a_{0}^{*}}\enspace\text{as well as}\enspace$
$\bigl{(}\mathbb{A}^{\mathrm{H}}(u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})-\bigl{(}B_{1}(u_{0}^{*})-B_{1}(u_{0})\bigr{)}\binom{\widehat{h}+h_{0}^{*}}{\widehat{a}+a_{0}^{*}}.$
By Subsection 5.2, the latter two terms admit estimates by $CR(R+C_{T}^{*})$ and $C\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma}}(R+C_{T}^{*})$, respectively, and we observe that $\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}$ tends to zero as $T\to 0$ by Section 4(c). It thus remains to estimate (5.12) in the last part of the proof.
We start with $(\tilde{\mathbb{A}}^{\mathrm{H}}(\widehat{u}+u_{0}^{*})-\mathbb{A}^{\mathrm{H}}(\widehat{u}+u_{0}^{*}))(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})$. Below, we concentrate on the principal part and remark that the remaining part can be handled in a similar way. Recalling the transformed Hibler operator $\tilde{\mathbb{A}}^{\mathrm{H}}$ from (3.7) and the original Hibler operator from (3.3), we note that we need to estimate
(5.13) $\sum_{j,k,l,m=1}^{2}\bigl{(}a_{ij}^{klm}(\tilde{\varepsilon}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*}),P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*}))\partial_{m}\tilde{\varepsilon}_{jl}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})$
$\quad-a_{ij}^{kl}(\varepsilon(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*}),P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*}))\partial_{k}\varepsilon_{jl}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})\bigr{)}=\mathrm{I}+\mathrm{II}$
in the sequel. To ease notation, we write $\varepsilon$ and $\tilde{\varepsilon}$ as well as $P$, since the latter quantities always depend on $\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*}$ and $\widehat{h}+h_{0}^{*}$ as well as $\widehat{a}+a_{0}^{*}$, respectively. The terms $\mathrm{I}$ and $\mathrm{II}$ in (5.13) are defined by
$\mathrm{I}\coloneqq\sum_{j,k,l,m=1}^{2}\bigl{(}a_{ij}^{klm}(\tilde{\varepsilon},P)\partial_{m}\tilde{\varepsilon}_{jl}-a_{ij}^{kl}(\varepsilon,P)\partial_{k}\tilde{\varepsilon}_{jl}\bigr{)}\enspace\text{and}\enspace\mathrm{II}\coloneqq\sum_{j,k,l=1}^{2}\bigl{(}a_{ij}^{kl}(\varepsilon,P)\partial_{k}\tilde{\varepsilon}_{jl}-a_{ij}^{kl}(\varepsilon,P)\partial_{k}\varepsilon_{jl}\bigr{)}.$
With regard to $\mathrm{II}$, it is necessary to estimate the difference of $\tilde{\varepsilon}$ and $\varepsilon$. In fact, we compute the difference $\partial_{k}\tilde{\varepsilon}_{jl}-\partial_{k}\varepsilon_{jl}$ to be given by
$\frac{1}{2}\sum_{n=1}^{2}\bigl{(}(\partial_{j}Y_{n}-\delta_{jn})\partial_{k}\partial_{n}\tilde{v}_{\mathrm{ice},l}+(\partial_{l}Y_{n}-\delta_{ln})\partial_{k}\partial_{n}\tilde{v}_{\mathrm{ice},j}+(\partial_{k}\partial_{j}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}+(\partial_{k}\partial_{l}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}\bigr{)}.$
From Subsection 5.2, we derive that
$\|(\partial_{j}Y_{n}-\delta_{jn})\partial_{k}\partial_{n}\tilde{v}_{\mathrm{ice},l}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}$ $\leq C\cdot\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigl{(}\|\widehat{u}\|_{\mathbb{E}_{1}}+\|u_{0}^{*}\|_{\mathbb{E}_{1}}\bigr{)}$
$\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}(R+C_{T}^{*}).$
The term $(\partial_{l}Y_{n}-\delta_{ln})\partial_{k}\partial_{n}\tilde{v}_{\mathrm{ice},j}$ can be treated in the same way. Employing the embedding of the trace space $\mathrm{X}_{\gamma}$ from (2.7), Subsection 5.2(c), Subsection 5.2, the trace space embedding from Subsection 5.2(a) with $T$-independent embedding constant and the estimate (5.11), we argue that
$\|(\partial_{k}\partial_{j}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}$ $\leq\|(\partial_{k}\partial_{j}Y_{n})\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\cdot\|\partial_{n}\tilde{v}_{\mathrm{ice},j}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}$
$\leq C\cdot\|\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigl{(}\|\widehat{v}_{\mathrm{ice}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}^{1}(\overline{\Omega}))}+\|v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{C}^{1}(\overline{\Omega}))}\bigr{)}$
$\leq CT^{\nicefrac{{1}}{{p}}}\cdot\|\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigl{(}\|\widehat{u}\|_{\mathrm{BUC}(0,T;\mathrm{X}_{\gamma})}+\|u_{0}^{*}\|_{\mathrm{BUC}(0,T;\mathrm{X}_{\gamma})}\bigr{)}$
$\leq CT^{\nicefrac{{1}}{{p}}}(R+C_{0}+C_{T}^{*}).$
As $(\partial_{k}\partial_{l}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}$ can be handled likewise, we have an estimate of the difference $\partial_{k}\tilde{\varepsilon}_{jl}-\partial_{k}\varepsilon_{jl}$. The shape of the coefficients as introduced in (3.4), $\tilde{u}(t)\in V$ for all $t\in[0,T]$ as stated in Subsection 5.1 and the estimate of $\tilde{u}$ from (5.6) reveal the boundedness, i. e., $\|a_{ij}^{kl}(\varepsilon,P)\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\leq C$. A concatenation of the previous estimates yields the existence of $\beta>0$ such that
$\|\mathrm{II}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\beta}(R+C_{0}+C_{T}^{*}).$
Invoking $a_{ij}^{klm}$ from (3.6), we further decompose $\mathrm{I}$ into $\mathrm{III}$ and $\mathrm{IV}$ given by
$\mathrm{III}\coloneqq\sum_{j,k,l,m=1}^{2}\bigl{(}a_{ij}^{klm}(\tilde{\varepsilon},P)-a_{ij}^{klm}(\varepsilon,P)\bigr{)}\partial_{m}\tilde{\varepsilon}_{jl}\enspace\text{and}\enspace\mathrm{IV}\coloneqq\sum_{j,k,l,m=1}^{2}a_{ij}^{kl}(\partial_{k}Y_{m}-\delta_{km})\partial_{m}\tilde{\varepsilon}_{jl}.$
With regard to $\mathrm{IV}$, we conclude from Subsection 5.2 that
(5.14) $\|\partial_{k}Y_{m}-\delta_{km}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\leq C\cdot\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}.$
Concerning $\partial_{m}\tilde{\varepsilon}_{jl}$, we can proceed in a similar way as above to derive $\|\partial_{m}\tilde{\varepsilon}_{jl}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq C$. Together with the above boundedness of the coefficients, this leads to
(5.15) $\|\mathrm{IV}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}.$
The last term in $\mathrm{III}$ can be estimated as described above. In addition, $\partial_{j}Y_{k}$ is bounded from above by a constant in $\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))$ in view of Subsection 5.2. Consequently, the task boils down to an estimate of the difference $a_{ij}^{kl}(\tilde{\varepsilon},P)-a_{ij}^{kl}(\varepsilon,P)$. The difference only concerns $\varepsilon$ and $\tilde{\varepsilon}$, so the first factor can be bounded similarly as above. Thanks to the smooth dependence on $\varepsilon$ in view of the regularization by $\delta>0$, we derive from the mean value theorem that
$\biggl{\|}\frac{1}{\triangle_{\delta}(\tilde{\varepsilon})}\left(\mathbb{S}_{ij}^{kl}-\frac{1}{\triangle_{\delta}^{2}(\tilde{\varepsilon})}(\mathbb{S}\tilde{\varepsilon})_{ik}(\mathbb{S}\tilde{\varepsilon})_{jl}\right)-\frac{1}{\triangle_{\delta}(\varepsilon)}\left(\mathbb{S}_{ij}^{kl}-\frac{1}{\triangle_{\delta}^{2}(\varepsilon)}(\mathbb{S}\varepsilon)_{ik}(\mathbb{S}\varepsilon)_{jl}\right)\biggr{\|}_{\infty,\infty}\leq C\cdot\|\tilde{\varepsilon}-\varepsilon\|_{\infty,\infty},$
where $\|\cdot\|_{\infty,\infty}=\|\cdot\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}$. Moreover, we compute
$\tilde{\varepsilon}_{ij}-\varepsilon_{ij}=\frac{1}{2}\sum_{k=1}^{2}\bigl{(}((\partial_{i}Y_{k})-\delta_{ik})\partial_{k}\tilde{v}_{\mathrm{ice},j}+((\partial_{j}Y_{k})-\delta_{jk})\partial_{k}\tilde{v}_{\mathrm{ice},i}\bigr{)}.$
The differences $(\partial_{i}Y_{k})-\delta_{ik}$ and $(\partial_{j}Y_{k})-\delta_{jk}$ can be handled with (5.14) to obtain an estimate by a $T$-power. Concerning the derivative of $\tilde{v}_{\mathrm{ice}}$, we employ (2.7), Subsection 5.2(a) and (5.11) to get
$\|\partial_{k}\tilde{v}_{\mathrm{ice}}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\leq C\bigl{(}\|\widehat{u}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}+\|u_{0}^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\bigr{)}\leq C(R+C_{0}+C_{T}^{*}).$
This yields that
(5.16) $\|\mathrm{III}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}(R+C_{0}+C_{T}^{*}).$
Putting together (5.15) and (5.16), we conclude that
$\|\mathrm{I}\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}(1+R+C_{0}+C_{T}^{*}).$
Concerning the corresponding off-diagonal part $(\tilde{B}(\widehat{u}+u_{0}^{*})-B_{1}(\widehat{u}+u_{0}^{*}))\binom{\widehat{h}+h_{0}^{*}}{\widehat{a}+a_{0}^{*}}$, we calculate the $h$-part
$\frac{\partial_{h}P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*})}{2\rho_{\mathrm{ice}}(\widehat{h}+h_{0}^{*})}\sum_{j=1}^{2}(\partial_{i}Y_{j}-\delta_{ij})\partial_{j}(\widehat{h}+h_{0}^{*}).$
Again, for the inverse of $2\rho_{\mathrm{ice}}(\widehat{h}+h_{0}^{*})$, we can use (5.10), while $\partial_{h}P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*})=p^{*}\mathrm{e}^{-c_{\bullet}(1-(\widehat{a}+a_{0}^{*}))}$. Upon deducing from Subsection 5.1 that $\tilde{a}(t)=\widehat{a}(t)+a_{0}^{*}(t)\in(0,1)$ for all $t\in[0,T]$, we find that
$\|\partial_{h}P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*})\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\leq C.$
In total, additionally exploiting the estimate of $\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y$ from Subsection 5.2 and Subsection 5.2(c), we obtain
$\quad\biggl{\|}\frac{\partial_{h}P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*})}{2\rho_{\mathrm{ice}}(\widehat{h}+h_{0}^{*})}\sum_{j=1}^{2}(\partial_{i}Y_{j}-\delta_{ij})\partial_{j}(\widehat{h}+h_{0}^{*})\biggr{\|}_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}$
$\leq C\cdot\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{L}^{\infty}(\Omega))}\bigl{(}\|\widehat{h}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|h_{0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigr{)}$
$\leq C\cdot\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigl{(}T^{\nicefrac{{1}}{{p}}}\cdot\|\widehat{h}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|u_{0}^{*}\|_{\mathbb{E}_{1}}\bigr{)}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}\bigl{(}T^{\nicefrac{{1}}{{p}}}R+C_{T}^{*}\bigr{)}.$
Analogously, we derive an estimate of the term $\frac{\partial_{a}P(\widehat{h}+h_{0}^{*},\widehat{a}+a_{0}^{*})}{2\rho_{\mathrm{ice}}(\widehat{h}+h_{0}^{*})}\sum_{j=1}^{2}(\partial_{i}Y_{j}-\delta_{ij})\partial_{j}(\widehat{a}+a_{0}^{*})$.
The assertion of the proposition then follows for some $C_{F_{1}}(R,T)>0$. A choice of $R>0$ sufficiently small and then $T\to 0$ imply $C_{F_{1}}(R,T)<\nicefrac{{R}}{{6C_{\mathrm{MR}}}}$ in view of Section 4(b) and (c), where the latter ensure $C_{T}^{*}\to 0$ and $\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$. ∎
The next terms to be treated are $F_{2}$ and $F_{3}$.
[…]
Let $p,q\in(1,\infty)$ such that (2.6) is valid, consider $\tilde{u}\coloneqq\widehat{u}+u_{0}^{*}$, where $\widehat{u}\in\prescript{}{0}{\mathbb{E}_{1}}$ and $\|\widehat{u}\|_{\mathbb{E}_{1}}\leq R$, and $u_{0}^{*}$ represents the reference solution from Section 4. Moreover, assume that $f_{\mathrm{gr}}$ fulfills Section 2(c). Then for the $T$-independent maximal regularity constant $C_{\mathrm{MR}}>0$ from Section 4, there are $C_{F_{2}}(R,T)$, $C_{F_{3}}(R,T)>0$ with $C_{F_{2}}(R,T)$, $C_{F_{3}}(R,T)<\nicefrac{{R}}{{6C_{\mathrm{MR}}}}$, for $R>0$ and $T>0$ sufficiently small, and $\|F_{2}(\widehat{u})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq C_{F_{2}}(R,T)$ and $\|F_{3}(\widehat{u})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq C_{F_{3}}(R,T)$.
Proof.
Conceptually, the proof parallels the one of Subsection 5.2, and we often exploit the Banach algebra structure of $\mathrm{W}^{1,q}(\Omega)$. Thanks to $\omega>0$ being fixed and Subsection 5.2(c) and (a), we first obtain
$\|\omega\tilde{h}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq C\bigl{(}T^{\nicefrac{{1}}{{p}}}\cdot\|\widehat{u}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}+C_{T}^{*}\bigr{)}\leq C\bigl{(}T^{\nicefrac{{1}}{{p}}}R+C_{T}^{*}\bigr{)}.$
With regard to Section 2(c) as well as Subsection 5.2(c) and (a), we then get
$\|S_{\mathrm{h}}(\tilde{h},\tilde{a})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq\biggl{\|}f_{\mathrm{gr}}\biggl{(}\frac{\tilde{h}}{\tilde{a}}\biggr{)}\biggr{\|}_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\cdot\|\widehat{a}+a_{0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$
$\quad+\|f_{\mathrm{gr}}(0)\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|f_{\mathrm{gr}}(0)(\widehat{a}+a_{0}^{*})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$
$\leq C\bigl{(}T^{\nicefrac{{1}}{{p}}}+\|\widehat{a}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}+\|a_{0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\bigr{)}$
$\leq C\bigl{(}T^{\nicefrac{{1}}{{p}}}(1+\|\widehat{u}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})})+C_{T}^{*}\bigr{)}\leq C\bigl{(}T^{\nicefrac{{1}}{{p}}}(1+R)+C_{T}^{*}\bigr{)}.$
Hence, it remains to estimate the term $h_{0}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}-\tilde{h}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}\tilde{v}_{\mathrm{ice},j}$. We treat this term by the addition and subtraction of the intermediate terms $(\widehat{h}+h_{0}^{*})\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})$ as well as $h_{0}^{*}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})$. In that respect, we get
$\|(h_{0}-h_{0}^{*})\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\cdot\|h_{0}-h_{0}^{*}\|_{\mathrm{BUC}([0,T];\mathrm{W}^{1,q}(\Omega))}\|\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega))}$
$\leq C\cdot\|u_{0}-u_{0}^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}(R+C_{T}^{*}).$
Concerning $h_{0}^{*}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})-(\widehat{h}+h_{0}^{*})\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})=\widehat{h}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})$, Subsection 5.2(a) yields
$\|\widehat{h}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$ $\leq C\cdot\|\widehat{h}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}(\|\widehat{v}_{\mathrm{ice}}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega))}+\|v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega))})$
$\leq C\cdot\|\widehat{u}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\bigl{(}\|\widehat{u}\|_{\mathbb{E}_{1}}+\|u_{0}^{*}\|_{\mathbb{E}_{1}}\bigr{)}\leq CR(R+C_{T}^{*}).$
The last term in the context of $F_{2}$ can be written as
$\tilde{h}\mathrm{div}_{\mathrm{H}}\,\tilde{v}_{\mathrm{ice}}-\tilde{h}\sum_{j,k=1}^{2}(\partial_{j}Y_{k})\partial_{k}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})_{j}=(\widehat{h}+h_{0}^{*})\sum_{j,k=1}^{2}(\delta_{jk}-\partial_{j}Y_{k})\partial_{k}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})_{j}.$
For its treatment, we make use of Subsection 5.2, the shape of $\mathrm{X}_{\gamma}$ as indicated in (2.5), Subsection 5.2(a) and (5.11) to derive
$\quad\left\|(\widehat{h}+h_{0}^{*})\sum_{j,k=1}^{2}(\delta_{jk}-\partial_{j}Y_{k})\partial_{k}(\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*})_{j}\right\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}$
$\leq C\cdot\|\widehat{h}+h_{0}^{*}\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\cdot\|\operatorname{Id}_{2}-\nabla_{\mathrm{H}}Y\|_{\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,q}(\Omega))}\cdot\|\widehat{v}_{\mathrm{ice}}+v_{\mathrm{ice},0}^{*}\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{2,q}(\Omega))}$
$\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}\bigl{(}\|\widehat{u}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}+\|u_{0}^{*}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\bigr{)}\bigl{(}\|\widehat{u}\|_{\mathbb{E}_{1}}+\|u_{0}^{*}\|_{\mathbb{E}_{1}}\bigr{)}\leq CT^{\nicefrac{{1}}{{p^{\prime}}}}(R+C_{0}+C_{T}^{*})(R+C_{T}^{*}).$
The assertion of the proposition follows from a combination of the previous estimates and the observation that $C_{T}^{*}\to 0$ and $\|u_{0}^{*}-u_{0}\|_{\mathrm{BUC}([0,T];\mathrm{X}_{\gamma})}\to 0$ as $T\to 0$ by Section 4(b) and (c). ∎
For completeness, we provide the differences of $F_{1}$, $F_{2}$ and $F_{3}$. In fact, for $\widehat{u}_{i}=\tilde{u}_{i}-u_{0}^{*}$, $i=1,2$, we get
$F_{1}(\widehat{u}_{1})-F_{1}(\widehat{u}_{2})=\bigl{(}\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u}_{1})-\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u}_{2})\bigr{)}\tilde{v}_{\mathrm{ice},1}-\bigl{(}\tilde{B}(\tilde{u}_{1})-\tilde{B}(\tilde{u}_{2})\bigr{)}\binom{\tilde{h}_{1}}{\tilde{a}_{1}}$
$\quad+\bigl{(}\tilde{\mathbb{A}}^{\mathrm{H}}(\tilde{u}_{2})-\mathbb{A}^{\mathrm{H}}(u_{0})\bigr{)}(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})-\bigl{(}\tilde{B}(\tilde{u}_{2})-B_{1}(u_{0})\bigr{)}\binom{\widehat{h}_{1}-\widehat{h}_{2}}{\widehat{a}_{1}-\widehat{a}_{2}}+\omega(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})$
$\quad-c_{\mathrm{cor}}(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})^{\perp}+\biggl{(}\frac{1}{\rho_{\mathrm{ice}}\tilde{h}_{1}}-\frac{1}{\rho_{\mathrm{ice}}\tilde{h}_{2}}\biggr{)}\bigl{(}\tau_{\mathrm{atm}}+\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice},1})\bigr{)}+\frac{1}{\rho_{\mathrm{ice}}\tilde{h}_{2}}\bigl{(}\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice},1})-\tau_{\mathrm{ocn}}(\tilde{v}_{\mathrm{ice},2})\bigr{)},$
$F_{2}(\widehat{u}_{1})-F_{2}(\widehat{u}_{2})$ $=-\sum_{j,k=1}^{2}\bigl{[}\tilde{h}_{1}(\partial_{j}Y_{k}^{(1)})-\tilde{h}_{2}(\partial_{j}Y_{k}^{(2)})\bigr{]}\partial_{k}(\tilde{v}_{\mathrm{ice},1})_{j}-\biggl{[}\tilde{h}_{2}\sum_{j,k=1}^{2}(\partial_{j}Y_{k}^{(2)})\partial_{k}(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})_{j}$
$\qquad-h_{0}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})\biggr{]}+\omega(\widehat{h}_{1}-\widehat{h}_{2})+S_{\mathrm{h}}(\tilde{h}_{1},\tilde{a}_{1})-S_{\mathrm{h}}(\tilde{h}_{2},\tilde{a}_{2})\enspace\text{and}\enspace$
$F_{3}(\widehat{u}_{1})-F_{3}(\widehat{u}_{2})$ $=-\sum_{j,k=1}^{2}\bigl{[}\tilde{a}_{1}(\partial_{j}Y_{k}^{(1)})-\tilde{a}_{2}(\partial_{j}Y_{k}^{(2)})\bigr{]}\partial_{k}(\tilde{v}_{\mathrm{ice},1})_{j}-\biggl{[}\tilde{a}_{2}\sum_{j,k=1}^{2}(\partial_{j}Y_{k}^{(2)})\partial_{k}(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})_{j}$
$\qquad-a_{0}\mathrm{div}_{\mathrm{H}}\,(\widehat{v}_{\mathrm{ice},1}-\widehat{v}_{\mathrm{ice},2})\biggr{]}+\omega(\widehat{a}_{1}-\widehat{a}_{2})+S_{\mathrm{a}}(\tilde{h}_{1},\tilde{a}_{1})-S_{\mathrm{a}}(\tilde{h}_{2},\tilde{a}_{2}).$
The lemma below discusses the Lipschitz continuity of $F_{1}$ from (5.2). It can be obtained in a similar way as Subsection 5.2 when replacing the estimates from Subsection 5.2(a) by the (b)-part in order to estimate the difference in the diffeomorphism. We also refer here to [5, Lemma 6.4.8] for details.
[…]
Let $p,q\in(1,\infty)$ satisfy (2.6), and let $\tilde{u}_{i}\coloneqq\widehat{u}_{i}+u_{0}^{*}$, $i=1,2$, where $\widehat{u}_{i}\in\prescript{}{0}{\mathbb{E}_{1}}$ with $\|\widehat{u}_{i}\|_{\mathbb{E}_{1}}\leq R$, and $u_{0}^{*}$ denotes the reference solution from Section 4. In addition, suppose that $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}$, $\nabla_{\mathrm{H}}H$ and $f_{\mathrm{gr}}$ fulfill Section 2. Then there exists $L_{F_{1}}(R,T)>0$ with $L_{F_{1}}(R,T)\to 0$ as $R\to 0$ and $T\to 0$ as well as $\|F_{1}(\widehat{u}_{1})-F_{1}(\widehat{u}_{2})\|_{\mathrm{L}^{p}(0,T;\mathrm{L}^{q}(\Omega))}\leq L_{F_{1}}(R,T)\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}$.
Next, we state the Lipschitz estimates of the terms $F_{2}$ and $F_{3}$. They can also be derived in a similar manner as Subsection 5.2, see also [5, Lemma 6.4.9].
[…]
Let $p,q\in(1,\infty)$ fulfill (2.6), and consider $\tilde{u}_{i}\coloneqq\widehat{u}_{i}+u_{0}^{*}$, $i=1,2$, with $\widehat{u}_{i}\in\prescript{}{0}{\mathbb{E}_{1}}$ and $\|\widehat{u}_{i}\|_{\mathbb{E}_{1}}\leq R$, and $u_{0}^{*}$ represents the reference solution from Section 4. Moreover, assume that $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}$, $\nabla_{\mathrm{H}}H$ and $f_{\mathrm{gr}}$ satisfy Section 2. Then there are $L_{F_{2}}(R,T)>0$ and $L_{F_{3}}(R,T)>0$ with $L_{F_{j}}(R,T)\to 0$, $j=2,3$, as $R\to 0$ and $T\to 0$, $\|F_{2}(\widehat{u}_{1})-F_{2}(\widehat{u}_{2})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq L_{F_{2}}(R,T)\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}$ as well as $\|F_{3}(\widehat{u}_{1})-F_{3}(\widehat{u}_{2})\|_{\mathrm{L}^{p}(0,T;\mathrm{W}^{1,q}(\Omega))}\leq L_{F_{3}}(R,T)\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}$.
5.3. Proof of Theorem 3.1 and Theorem 2.1
This final subsection presents the proofs of the main theorem in Lagrangian and Eulerian coordinates.
Proof of Theorem 3.1.
We recall $\mathcal{K}_{T}^{R}$ and $\Phi_{T}^{R}$ from (5.4). For $\widehat{u}\in\mathcal{K}_{T}^{R}$, the maximal regularity from Section 4 and the self-map estimates from Subsection 5.2 and Subsection 5.2 imply that
$\|\Phi_{T}^{R}(\widehat{u})\|_{\mathbb{E}_{1}}\leq C_{\mathrm{MR}}\cdot\|(F_{1}(\widehat{u}),F_{2}(\widehat{u}),F_{3}(\widehat{u}))\|_{\mathbb{E}_{0}}\leq C_{\mathrm{MR}}\left(\frac{R}{6C_{\mathrm{MR}}}+\frac{R}{6C_{\mathrm{MR}}}+\frac{R}{6C_{\mathrm{MR}}}\right)\leq\frac{R}{2}$
for $R>0$ and $T>0$ sufficiently small by virtue of the $T$-independence of $C_{\mathrm{MR}}>0$. Furthermore, from Subsection 5.2 and Subsection 5.2, we deduce the existence of a Lipschitz constant $L(R,T)>0$ such that $L(R,T)\to 0$ as $R\to 0$ and $T\to 0$, and
$\|\Phi_{T}^{R}(\widehat{u}_{1})-\Phi_{T}^{R}(\widehat{u}_{2})\|_{\mathbb{E}_{1}}\leq C_{\mathrm{MR}}L(R,T)\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}\leq\frac{1}{2}\cdot\|\widehat{u}_{1}-\widehat{u}_{2}\|_{\mathbb{E}_{1}}$
for $\widehat{u}_{1}$, $\widehat{u}_{2}\in\mathcal{K}_{T}^{R}$ as well as $R>0$ and $T>0$ sufficiently small. In total, $\Phi_{T}^{R}$ is a self-map and contraction on $\mathcal{K}_{T}^{R}$, yielding the existence of a unique fixed point $u_{*}\in\prescript{}{0}{\mathbb{E}_{1}}$ of (5.1). In other words, (3.8) admits a unique solution $u_{*}\coloneqq\widehat{u}+u_{0}^{*}$. Together with Subsection 5.2(a), we obtain
$\tilde{u}\in\mathbb{E}_{1}=\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1})\hookrightarrow\mathrm{BUC}([0,T];\mathrm{X}_{\gamma}).$
Finally, it is a consequence of Subsection 5.1 that $\tilde{u}(t)\in V$ for all $t\in[0,T]$, completing the proof. ∎
The last step is to deduce Theorem 2.1 from Theorem 3.1.
Proof of Theorem 2.1.
Because of $T\leq T_{1}$, we argue that $X(t,\cdot)$ is in particular invertible on $\overline{\Omega}$ with inverse $Y(t,\cdot)=[X(t,\cdot)]^{-1}$. From the representation of $X$ in (5.5) as well as the regularity of $\tilde{v}_{\mathrm{ice}}$ in Theorem 3.1, it follows that $X\in\mathrm{W}^{2,p}(0,T;\mathrm{L}^{q}(\Omega)^{2})\cap\mathrm{W}^{1,p}(0,T;\mathrm{W}^{2,q}(\Omega)^{2})$. This is also valid for the inverse $Y$. As a result, the variables in Eulerian coordinates $u=(v_{\mathrm{ice}},h,a)$ can be recovered from those in Lagrangian coordinates by $v_{\mathrm{ice}}(t,x_{\mathrm{H}})\coloneqq\tilde{v}_{\mathrm{ice}}(t,Y(t,x_{\mathrm{H}}))$, $h(t,x_{\mathrm{H}})\coloneqq\tilde{h}(t,Y(t,x_{\mathrm{H}}))$ as well as $a(t,y_{\mathrm{H}})\coloneqq\tilde{a}(t,Y(t,x_{\mathrm{H}}))$. The proof is completed upon observing that the change of coordinates does not affect the regularity properties, and it has no influence on the property $u\in\mathrm{C}([0,T];V)$ either, while the uniqueness is implied by the uniqueness of the fixed point joint with the uniqueness of the transform. ∎
We conclude this section by proving the continuous dependence on the initial data as asserted in Section 3 as well as the blow-up criterion from Section 3.
Proof of Section 3.
This result can be obtained by a slight modification of the fixed point argument. Indeed, for $u_{0}^{*}$ denoting the reference solution associated to initial value $u_{0}$ and resulting from Section 4 as well as $T>0$ and $R>0$, we set $\mathcal{K}_{T,u_{i}}^{R}\coloneqq\left\{\overline{u}\in\mathbb{E}_{1}:\left.\overline{u}\right|_{t=0}=u_{i}\enspace\text{and}\enspace\|\overline{u}-u_{0}^{*}\|_{\mathbb{E}_{1}}\leq r\right\}$, $i=1,2$. Moreover, we define $\Phi_{T,u_{i}}^{R}\colon\mathcal{K}_{T,u_{i}}^{R}\to\mathbb{E}_{1}$, $\Phi_{T,u_{i}}^{R}(\overline{u})=\widehat{u}$ to be the solution to the linearized problem as in (5.1), with right-hand sides $F_{1}(\overline{u})$, $F_{2}(\overline{u})$ and $F_{3}(\overline{u})$ as introduced in (5.2), and with initial values $u_{i}=(v_{\mathrm{ice},i},h_{i},a_{i})$. Following the proof of [17, Theorem 2.1], and adapting the arguments from the nonlinear estimates in Subsection 5.2, for $R>0$, $T>0$ and $r>0$ sufficiently small, we find that
$\|\Phi_{T,u_{1}}^{R}(\overline{u}_{1})-\Phi_{T,u_{2}}^{R}(\overline{u}_{2})\|_{\mathbb{E}_{1}}\leq\frac{1}{2}\cdot\|\overline{u}_{1}-\overline{u}_{2}\|_{\mathbb{E}_{1}}+C\cdot\|u_{1}-u_{2}\|_{\mathrm{X}_{\gamma}}$
for all $u_{1}$, $u_{2}\in\overline{\mathbb{B}}_{\mathrm{X}_{\gamma}}(u_{0},r)$ as well as $\overline{u}_{1}\in\mathcal{K}_{T,u_{1}}^{R}$ and $\overline{u}_{2}\in\mathcal{K}_{T,u_{2}}^{R}$. Applying the preceding estimate to the solutions $u(\cdot,u_{1})$, $u(\cdot,u_{2})$, whose existence can be obtained in a similar manner as in the proof of Theorem 3.1 upon slightly modifying the arguments, we deduce the assertion of the corollary. ∎
The second result follows similarly as [30, Corollary 5.1.2]. For convenience, we provide a short proof.
Proof of Section 3.
[…]
6. Concluding remarks
In this final section, we discuss possible extensions of the results presented in the previous sections, sketch possible difficulties and mention remaining problems.
In order to fully exploit the parabolic regularization, it would be possible to introduce time weights of the shape $t^{1-\mu}$ for $\mu\in(\nicefrac{{1}}{{p}},1]$ and to consider $\mathrm{L}_{\mu}^{p}(0,T;\mathrm{X}_{0})$ instead of $\mathrm{L}^{p}(0,T;\mathrm{X}_{0})$, where $f\in\mathrm{L}_{\mu}^{p}(0,T;\mathrm{X}_{0})$ if and only if $t^{1-\mu}f\in\mathrm{L}^{p}(0,T;\mathrm{X}_{0})$. The spaces $\mathrm{W}_{\mu}^{1,p}(0,T;\mathrm{X}_{0})$ and $\mathrm{L}_{\mu}^{p}(0,T;\mathrm{X}_{1})$ are defined analogously. The underlying data space then is $\mathbb{E}_{0,\mu}\coloneqq\mathrm{L}_{\mu}^{p}(0,T;\mathrm{X}_{0})$, while the associated time trace space becomes $\mathrm{X}_{\gamma,\mu}\coloneqq(\mathrm{X}_{0},\mathrm{X}_{1})_{\mu-\nicefrac{{1}}{{p}},p}$, and the corresponding maximal regularity space reads as
$\mathbb{E}_{1,\mu}\coloneqq\mathrm{W}_{\mu}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}_{\mu}^{p}(0,T;\mathrm{X}_{1}).$
An implementation of time weights would require an adjustment of the nonlinear estimates. For the theory of critical spaces in the context of evolution equations and weighted spaces, we also refer to the work of Prüss, Simonett and Wilke [31].
The bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix $A(u_{1})$ from (4.3) as proved in Theorem 4.1 paves the way for interesting applications such as the theory of Banach scales due to Amann [2, Chapter 5] or stochastic maximal regularity, see also [27] for more details on this property.
The investigation of global-in-time results as e. g. obtained in [6, Theorem 2.3], see also [5, Theorem 3.6.6], is left to future study for the present problem. In fact, there are several difficulties here. First, the linearized operator matrix $A(u_{*})$ at a constant equilibrium solution of the shape $u_{*}=(0,h_{*},a_{*})\in V$ seems to be “more singular” than in the fully parabolic situation. Indeed, zero is no longer a semi-simple eigenvalue, and the lack of invertibility cannot be overcome as for instance in [5, Section 7.2] by restricting to functions with spatial average zero. Another difficulty arises in the transfer of global-in-time results from the Lagrangian to the Eulerian setting. For this, it is necessary to have a global-in-time Lagrangian change of coordinates at hand. In [10], Danchin, Hieber, Mucha and Tolksdorf address this aspect in the context of free boundary value problems in the half space by using Da Prato-Grisvard theory, allowing for maximal $\mathrm{L}^{1}$-regularity in the critical (homogeneous) Besov space. For a different approach to such problems, we also refer to the work of Ogawa and Shimizu [28].
One can also ask whether the results established in [7], [3], [4] remain valid in the present situation of the parabolic-hyperbolic problem. We also leave this to future study.
Acknowledgements The author would like to thank Tim Binz, Karoline Disser and Matthias Hieber for fruitful discussions. Moreover, he acknowledges the support by DFG project FOR 5528. Most of the results in this paper are contained in modified form in Chapter 6 of the PhD thesis [5] of the author.

### v3

@@
 It is the aim of the present paper to prove the local strong well-posedness of the parabolic-hyperbolic variant of Hibler’s model, where the most commonly used regularization for the stress tensor is employed.
-Unlike the models under consideration in [6] or [23], this is precisely the model typically studied in numerical simulations.
+Unlike the models considered in [6] or [23], this is the model typically studied in numerical simulations.
 In contrast to the rigorous mathematical analysis, which only started quite recently, there is vast literature on numerical analysis and simulations of Hibler’s model, see for example the work of Zhang and Hibler [37].
@@
 Let us also mention the recent work of Piersanti and Temam [29] on the modeling and analysis of the dynamics of shallow ice sheets.
+For convenience of the reader, we briefly elaborate on the function spaces used in this paper.
+Consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with boundary of class $\mathrm{C}^{2}$, and suppose that $G$ either represents $\Omega$, or an interval of the form $(0,T)$, for $T>0$.
+With $p\in[1,\infty]$ and $k\in\mathbb{N}$, we denote by $\mathrm{L}^{p}(G)$ and $\mathrm{W}^{k,p}(G)$ the Lebesgue and Sobolev spaces on $G$, respectively.
+The space $\mathrm{W}_{0}^{1,q}(\Omega)$, $q\in(1,\infty)$, contains the functions in $\mathrm{W}^{1,q}(\Omega)$ whose trace vanishes on $\partial\Omega$.
+Moreover, for $l\in\{0,1\}$ and $\alpha\in(0,1)$, the space $\mathrm{C}^{l,\alpha}(\overline{\Omega})$ consists of all functions $f\in\mathrm{C}^{l}(\overline{\Omega})$ such that all partial derivatives $\partial^{\beta}f$ of orders $|\beta|\leq l$ are $\alpha$-Hölder continuous on $\overline{\Omega}$.
+By $\mathrm{C}_{\mathrm{b}}^{1}([0,\infty))$, we denote the functions in $\mathrm{C}^{1}([0,\infty))$ that are bounded on $[0,\infty)$, and whose derivative is additionally bounded on $[0,\infty)$.
+The space $\mathrm{BUC}([0,T])$ represents the bounded and uniformly continuous functions on $[0,T]$.
+For $s>0$ and $p$, $q\in(1,\infty)$, we use $\mathrm{B}_{qp}^{s}(\Omega)$ for the Besov spaces on $\Omega$.
+Note that they can be obtained from the Lebesgue and Sobolev spaces by real interpolation, see e.
+g.
+[35, Sec.
+2.4 and 4.3.1].
+More precisely, for $s$, $p$, $q$ as above, we have $\mathrm{B}_{qp}^{s}(\Omega)=(\mathrm{L}^{q}(\Omega),\mathrm{W}^{2,q}(\Omega))_{\frac{s}{2},p}$.
+Finally, for a Banach space $\mathrm{X}$, we denote by $\mathrm{L}^{p}(0,T;\mathrm{X})$ and $\mathrm{W}^{1,p}(0,T;\mathrm{X})$ the respective Bochner spaces.
 This article is organized as follows.
@@
 (2.1) $P=P(h,a)=p^{*}h\mathrm{e}^{-c_{\bullet}(1-a)}.$
-For the ratio $e>1$ of the major to minor axes of the elliptical yield curve of the principle components of the stress, we introduce
+It can be found in [13, p. 822, eq. (17)]. For the ratio $e>1$ of the major to minor axes of the elliptical yield curve of the principle components of the stress, we introduce
 $\triangle^{2}(\varepsilon)\coloneqq\left(\varepsilon_{11}^{2}+\varepsilon_{22}^{2}\right)\Bigl{(}1+\frac{1}{e^{2}}\Bigr{)}+\frac{4}{e^{2}}\varepsilon_{12}^{2}+2\varepsilon_{11}\varepsilon_{22}\Bigl{(}1-\frac{1}{e^{2}}\Bigr{)}.$
-The resulting bulk and shear viscosities then take the shape
+The bulk and shear viscosities are as made precise in [13, p. 819, eq. (7) and (8)], and they take the shape
 $\zeta(\varepsilon,P)=\frac{P(h,a)}{2\triangle(\varepsilon)}\enspace\text{and}\enspace\eta(\varepsilon,P)=e^{-2}\zeta(\varepsilon,P).$
@@
 $\tau_{\mathrm{atm}}=\rho_{\mathrm{atm}}C_{\mathrm{atm}}|V_{\mathrm{atm}}|R_{\mathrm{atm}}V_{\mathrm{atm}}\enspace\text{and}\enspace\tau_{\mathrm{ocn}}(v_{\mathrm{ice}})=\rho_{\mathrm{ocn}}C_{\mathrm{ocn}}|V_{\mathrm{ocn}}-v_{\mathrm{ice}}|R_{\mathrm{ocn}}(V_{\mathrm{ocn}}-v_{\mathrm{ice}})$
-for densities $\rho_{\mathrm{atm}}>0$ and $\rho_{\mathrm{ocn}}>0$, drag coefficients $C_{\mathrm{atm}}>0$ and $C_{\mathrm{ocn}}>0$, velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ of the atmosphere and the ocean and rotation matrices $R_{\mathrm{atm}}$ and $R_{\mathrm{ocn}}$.
+for constant densities $\rho_{\mathrm{atm}}>0$ and $\rho_{\mathrm{ocn}}>0$, constant drag coefficients $C_{\mathrm{atm}}>0$ and $C_{\mathrm{ocn}}>0$, velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ of the atmosphere and the ocean and rotation matrices $R_{\mathrm{atm}}$ and $R_{\mathrm{ocn}}$.
 The balance laws of the mean ice thickness $h$ and the ice concentration $a$ are given by
@@
 (2.4) $\mathrm{X}_{0}\coloneqq\mathrm{L}^{q}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)\enspace\text{and}\enspace\mathrm{X}_{1}\coloneqq\mathrm{W}^{2,q}(\Omega)^{2}\cap\mathrm{W}_{0}^{1,q}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega),$
-where $q\in(1,\infty)$.
-With regard to the initial values, for $p\in(1,\infty)$, we then introduce the time trace space $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$ taking the shape
+where $q\in(1,\infty)$. In the sequel, we will assume that $p$, $q\in(1,\infty)$ satisfy
+(2.5) $\frac{1}{p}+\frac{1}{q}<\frac{1}{2}.$
+With regard to the initial data, for $p$, $q\in(1,\infty)$ fulfilling (2.5), we then introduce the time trace space $\mathrm{X}_{\gamma}=(\mathrm{X}_{0},\mathrm{X}_{1})_{1-\nicefrac{{1}}{{p}},p}$ taking the shape
 (2.6) $\mathrm{X}_{\gamma}=\mathrm{B}_{qp,\mathrm{D}}^{2-\nicefrac{{2}}{{p}}}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega).$
-In the above, the subscript D indicates Dirichlet boundary conditions. If $p$, $q\in(1,\infty)$ satisfy
-(2.6) $\frac{1}{p}+\frac{1}{q}<\frac{1}{2},$
-then it follows from Sobolev embeddings, see for example [35, Theorem 4.6.1], that
+In the above, the subscript D indicates Dirichlet boundary conditions.
+For $p$, $q\in(1,\infty)$ such that (2.5) holds true, it follows from Sobolev embeddings, see for example [35, Theorem 4.6.1], that
 (2.7) $\mathrm{X}_{\gamma}\hookrightarrow\mathrm{B}_{qp}^{2-\nicefrac{{2}}{{p}}}(\Omega)^{2}\times\mathrm{W}^{1,q}(\Omega)\times\mathrm{W}^{1,q}(\Omega)\hookrightarrow\mathrm{C}^{1,\alpha}(\overline{\Omega})^{2}\times\mathrm{C}^{0,\alpha}(\overline{\Omega})\times\mathrm{C}^{0,\alpha}(\overline{\Omega}),$
@@
 […]
-Let $q\in(1,\infty)$, and let $T_{1}>0$ sufficiently large to be determined precisely later. We make the following assumptions on the data.
+Let $q\in(1,\infty)$. We make the following assumptions on the data.
   1. (a)
-    The wind velocity $V_{\mathrm{atm}}$ and the ocean velocity $V_{\mathrm{ocn}}$ satisfy $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{2q}(\Omega)^{2})$.
+    The wind and ocean velocities $V_{\mathrm{atm}}$ and $V_{\mathrm{ocn}}$ satisfy $V_{\mathrm{atm}}$, $V_{\mathrm{ocn}}\in\mathrm{L}^{\infty}(0,t;\mathrm{L}^{2q}(\Omega)^{2})$ for all $t>0$.
   2. (b)
-    For the sea surface dynamic height $H$, it is valid that $\nabla_{\mathrm{H}}H\in\mathrm{L}^{\infty}(0,T_{1};\mathrm{L}^{q}(\Omega)^{2})$.
+    For the sea surface dynamic height $H$, it is valid that $\nabla_{\mathrm{H}}H\in\mathrm{L}^{\infty}(0,t;\mathrm{L}^{q}(\Omega)^{2})$ for all $t>0$.
   3. (c)
     It holds that $f_{\mathrm{gr}}\in\mathrm{C}_{\mathrm{b}}^{1}([0,\infty))$ for the ice growth rate $f_{\mathrm{gr}}$.
+The above assumptions on the integrability and regularity of the data are relatively mild, and it seems that they are physically reasonable.
+For examples of ice growth rate functions $f_{\mathrm{gr}}$, which generally depend on the season, we refer for instance to the ones considered by Hibler [13, p.
+821].
 We are now in the position to formulate the main result of this article.
@@
 (3.6) $a_{ij}^{klm}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}),P(\tilde{h},\tilde{a}))\coloneqq(\partial_{k}Y_{m})a_{ij}^{kl}(\tilde{\varepsilon}(\tilde{v}_{\mathrm{ice}}),P(\tilde{h},\tilde{a})).$
-Besides, we compute
+Let us observe that the coefficients $a_{ij}^{klm}$ depend on the time $t$ via $\partial_{k}Y_{m}$. For simplicity of notation, we do not write this dependence explicitly. Besides, we compute
 $\partial_{m}\tilde{\varepsilon}_{jl}(v_{\mathrm{ice}})=\frac{1}{2}\sum_{n=1}^{2}\bigl{(}(\partial_{m}\partial_{j}Y_{n})\partial_{n}\tilde{v}_{ice,l}+(\partial_{j}Y_{n})\partial_{m}\partial_{n}\tilde{v}_{\mathrm{ice},l}+(\partial_{m}\partial_{l}Y_{n})\partial_{n}\tilde{v}_{\mathrm{ice},j}+(\partial_{l}Y_{n})\partial_{m}\partial_{n}\tilde{v}_{\mathrm{ice},j}\bigr{)}.$
@@
 $\tilde{u}\in\mathrm{W}^{1,p}(0,T;\mathrm{X}_{0})\cap\mathrm{L}^{p}(0,T;\mathrm{X}_{1})\cap\mathrm{C}([0,T];V).$
-Next, we study the continuous dependence of the solution on the initial data.
+Next, we study the continuous dependence of the solution on the initial data. As a by-product, we also obtain the local uniformity of the time of existence $T>0$ with respect to the initial data.
 […]
@@
 More precisely, we establish the boundedness of the $\mathcal{H}^{\infty}$-calculus of the operator matrix associated to the linearization of the parabolic-hyperbolic sea ice model.
-This result is of independent interest.
-As a corollary, we especially obtain the maximal $\mathrm{L}^{p}$-regularity for the fixed point argument.
+This result is of independent interest due to its interesting functional analytic consequences for the operator matrix such as a characterization of its fractional power domains, or stochastic maximal regularity.
+We refer here also to the discussion in Section 6.
+As a corollary, we especially obtain the maximal $\mathrm{L}^{p}$-regularity.
+Note that this (weaker) property is sufficient for the fixed point argument.
 First, we recall the $\mathrm{L}^{q}$-realization of the linearized Hibler operator as well as the bounded $\mathcal{H}^{\infty}$-calculus that it admits.
@@
 In the sequel, we fix $\omega>\omega_{0}$ for $\omega_{0}>0$ from Section 4.
-The maximal regularity constant $C_{\mathrm{MR}}>0$ can in particular be chosen independent of $T>0$ provided homogeneous initial values are considered.
+Note that the maximal regularity constant $C_{\mathrm{MR}}>0$ from Section 4 generally depends on $T$.
+However, it can be chosen independent of $T>0$ provided homogeneous initial values are considered.
 In order reduce the considerations to this situation, we now discuss the so-called reference solution capturing the initial values.
@@
 For the theory of critical spaces in the context of evolution equations and weighted spaces, we also refer to the work of Prüss, Simonett and Wilke [31].
-The bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix $A(u_{1})$ from (4.3) as proved in Theorem 4.1 paves the way for interesting applications such as the theory of Banach scales due to Amann [2, Chapter 5] or stochastic maximal regularity, see also [27] for more details on this property.
+The bounded $\mathcal{H}^{\infty}$-calculus of the operator matrix $A(u_{1})$ from (4.3) as proved in Theorem 4.1 paves the way for interesting applications such as the theory of Banach scales due to Amann [2, Chapter 5] or the consideration of the problem in the stochastic setting via stochastic maximal regularity, see also [27] for more details on this property.
 The investigation of global-in-time results as e.
@@
 One can also ask whether the results established in [7], [3], [4] remain valid in the present situation of the parabolic-hyperbolic problem. We also leave this to future study.
-Acknowledgements The author would like to thank Tim Binz, Karoline Disser and Matthias Hieber for fruitful discussions.
-Moreover, he acknowledges the support by DFG project FOR 5528.
+Acknowledgements.
+The author would like to thank Tim Binz, Karoline Disser and Matthias Hieber for fruitful discussions.
+Moreover, he acknowledges the support by the German National Academy of Sciences Leopoldina through the Leopoldina Fellowship Program with grant number LPDS 2024-07 and the support by DFG project FOR 5528.
 Most of the results in this paper are contained in modified form in Chapter 6 of the PhD thesis [5] of the author.
+Data Availability Statement. The manuscript has no associated data.